### v1

# Second Law of Thermodynamics without Einstein Relation

Abstract
Materials that are constantly driven out of thermodynamic equilibrium, such as active and living systems, typically violate the Einstein relation. This may arise from active contributions to particle fluctuations which are unrelated to the dissipative resistance of the surrounding medium. We show that in these cases the widely used relation between informatic entropy production and heat dissipation does not hold. Consequently, fluctuation relations for the mechanical work, such as the Jarzynski and Crooks theorems, are invalid. We relate the breaking of the correspondence between entropy production and heat dissipation to the breaking of the fluctuation-dissipation theorem. We propose a temperature-like variable which restores this correspondence and gives rise to a generalized second law of thermodynamics, whereby the dissipated heat is necessarily nonnegative and vanishes at equilibrium. The Clausius inequality, Carnot maximum efficiency theorem, and relation between the maximum extractable work and the change of free energy are recovered as well.
Classical thermodynamics is the central theory for the possible forms of energy transfer in materials. It was historically formulated for the purpose of estimating the work that can be extracted from heat engines, utilizing the principle that heat spontaneously flows from “hotter” to “colder” objects. This is one version of the second law of thermodynamics, which determines the direction of spontaneous processes. Comparing the efficiencies of different heat engines provides a definition of a consistent “hotness” scale which is a thermodynamic state function, i.e., the temperature $T$ [1, 2]. In light of its vast applicability and usefulness for characterizing the behaviors of equilibrium materials, there is an ongoing effort to extend the theory of thermodynamics to systems that are constantly driven out of equilibrium, for example, self-propelled particles [3, 4, 5], chemical reactions [6, 7], and living systems [8]. Among the complications associated with nonequilibrium systems is the inadequacy of temperature and other variables as thermodynamic state functions [2, 9, 10, 11, 12, 13].
One of the consequences of thermodynamic equilibrium, where the system’s configurations are Boltzmann-distributed, is the Einstein relation (ER) [14, 15],
$\mathbf{D}=k_{\mathrm{B}}T\boldsymbol{\mu}\,.$ (1)
The diffusivity $\mathbf{D}$ and mobility $\boldsymbol{\mu}$ characterize, respectively, the strength of thermal fluctuations and the dissipative attenuation of motion. In general they may be space-dependent matrices, reflecting translational and rotational symmetry breaking. At equilibrium the effects of fluctuations and dissipation exactly balance [15], leading to Eq. (1). However, in cases where the forces are not conservative [16, 9], or if the fluctuations arise from coarse-graining of fast athermal or active forces [17, 18, 11, 19, 20, 21, 22], there should not be a generic relation between fluctuations (diffusivity) and dissipation (mobility). Thus, more often than not, in far-from-equilibrium cases such as active biopolymer networks [23, 24, 25], bacterial swarms [26, 27], and chemical reactions out of equilibrium [28, 7], the ER does not hold. In cases where $\mathbf{D}$ and $\boldsymbol{\mu}$ are proportional to one another, one can use Eq. (1) to define an effective temperature [2, 11, 17, 29]. However, in the general case these two position-dependent matrices need not be related by a constant uniform scalar, and Eq. (1) is violated.
One of the theories addressing nonequilibrium systems is stochastic thermodynamics [30]. By considering Langevin equations or discrete Markov chains it has succeeded in establishing trajectory-wise energy conservation (leading to the first law of thermodynamics) [31, 32, 33] and proving a wealth of fluctuation relations (FRs) [34], such as the Jarzynski equality [35, 36] and Crooks relation [37]. The latter also yields the second law of thermodynamics for the average entropy changes and heat increments (with some individual trajectories violating it [38]). Yet, implicitly, all the aforementioned FRs and extended laws of thermodynamics have relied on the existence of a temperature, defined through the ER, to connect the microscopic statistics to thermodynamic energy, mechanical and chemical work, and calorimentric heat.
Several definitions of a generalized nonequilibrium temperature were proposed based on generalizations of the fluctuation-dissipation theorem (FDT) [39, 11, 40, 18]. Most of these definitions still assume a thermal bath that obeys the ER. In other cases the ER is required to keep the extended FRs connected to physical measurables. For example, the Hatano-Sasa relation [41], an extension of the Jarzynski equality, requires Markovian dynamics only. It therefore relaxes Jarzynski’s assumption of a steady-state Boltzmann distribution for a given value of the external protocol [36], thus relaxing also the ER requirement. Yet, the effective Hamiltonian and work used by this relation coincide with their actual, measurable counterparts only under those extra assumptions. Otherwise, they demand knowledge of the complete microstates’ steady-state distribution, which is usually inaccessible or requires simplistic models [42]. The generalized temperature proposed here, arising directly from thermodynamics, does not rely on those assumptions.
We have two aims: (i) to investigate the fate of existing nonequilibrium FRs and extensions to the second law of thermodynamics once the assumption of an ER-based temperature is removed; and (ii) to formulate a generalized, ER-independent temperature and derive the resulting alternative thermodynamic identities.
Overdamped stochastic energetics.—We consider an ensemble of $N$ interacting degrees of freedom (particles), subjected to a deterministic external protocol $\Lambda_{t}$, which is a prechosen time-varying system parameter. The time-dependent stochastic microstate is an $N$-component vector, $\mathbf{X}_{t}$. We couple the particles to an arbitrary medium, which is assumed to provide them with effective mobility $\boldsymbol{\mu}(\mathbf{x},\lambda)$ and diffusivity $\mathbf{D}(\mathbf{x},\lambda)$ tensors, such that the particles follow the overdamped Langevin equation [Eq. (2) below]; no other property of the surrounding medium is specified. The aforementioned $\mathbf{X}_{t}$ and $\Lambda_{t}$ denote a particular time-dependent realizations of $\mathbf{x}$ and $\lambda$. The particles are subjected to an additional external force $\mathbf{F}_{\mathrm{ext}}(\mathbf{x},\lambda)$, which may be nonconservative and is not included in the internal energy. We denote the total ($N$-component) force acting on the particles by $\mathbf{F}=\mathbf{F}_{\mathrm{ext}}-\boldsymbol{\nabla}H$, where $H(\mathbf{x},\lambda)$ is the system’s Hamiltonian. The system is driven out of equilibrium by applying the nonconservative force $\mathbf{F}_{\mathrm{ext}}$ and/or violating the ER [Eq. (1)]. Thus, the system’s steady-state distribution is not Boltzmann, nor can a temperature be conventionally identified.
The system follows the overdamped Langevin equation,
$\dot{\mathbf{X}}_{t}$ $=$ $\boldsymbol{\mu}(\mathbf{X}_{t},\Lambda_{t})\cdot\mathbf{F}(\mathbf{X}_{t},\Lambda_{t})+{\boldsymbol{\nabla}}\cdot\mathbf{D}(\mathbf{X}_{t},\Lambda_{t})$ (2)
$+\mathbf{b}(\mathbf{X}_{t},\Lambda_{t})\cdot\boldsymbol{\eta}_{t}\,,$
where $\mathbf{b}$ is the noise magnitute satisfying $\mathbf{D}=\mathbf{b}\cdot\mathbf{b}^{\mathrm{T}}/2$, $\boldsymbol{\eta}_{t}$ is the normalized thermal noise [$\langle\boldsymbol{\eta}_{t}\rangle=\mathbf{0}$, $\langle\boldsymbol{\eta}_{t}\boldsymbol{\eta}_{t^{\prime}}\rangle=\mathbf{I}\delta(t-t^{\prime})$ where $\mathbf{I}$ is the identity matrix], and the equation is written in the Itô convention [43]. The so-called “spurious drift” term ${\boldsymbol{\nabla}}\cdot\mathbf{D}$ ensures that the Boltzmann distribution is obtained at equilibrium, i.e., in systems with a well-defined temperature, Eq. (1), under the effect of conservative forces [44, 45]. The corresponding Fokker-Planck equation [46] for the microstates distribution $p(\mathbf{x},t)$ is
$\frac{\partial p(\mathbf{x},t)}{\partial t}$ $=$ $-{\boldsymbol{\nabla}}\cdot\mathbf{J}(\mathbf{x},t)\,,$ (3)
$\mathbf{J}$ $=$ $\boldsymbol{\mu}\cdot\mathbf{F}\,p-\mathbf{D}\cdot{\boldsymbol{\nabla}}p\,,$ (4)
where $\mathbf{J}$ is the probability flux.
We now follow Sekimoto’s approach [31, 32] to identify the internal energy, work, and heat for trajectories in generic Langevin systems. The time-derivative of the Hamiltonian is [43]
$\dot{H}_{t}={\boldsymbol{\nabla}}H\circ\dot{\mathbf{X}}_{t}+\frac{\partial H}{\partial\lambda}\dot{\Lambda}_{t}\,,$ (5)
where $\circ$ is the Stratonovich product [47]. The term $(\partial H/\partial\lambda){\text{d}}\Lambda_{t}$ is the external work necessary to change the system’s internal energy over time ${\text{d}}t$. The force $\mathbf{F}_{\mathrm{ext}}(\mathbf{X}_{t},\Lambda_{t})$ acting along a displacement ${\text{d}}\mathbf{X}_{t}$ during ${\text{d}}t$ performs an additional work, $\mathbf{F}_{\mathrm{ext}}\circ{\text{d}}\mathbf{X}_{t}$. Thus the total work rate is
$\dot{W}_{t}=\frac{\partial H}{\partial\lambda}\dot{\Lambda}_{t}+\mathbf{F}_{\mathrm{ext}}\circ\dot{\mathbf{X}}_{t}\,.$ (6)
According to the first law of thermodynamics, the heat rate at time $t$ is the remaining energy flow into the system,
$\dot{Q}_{t}=\dot{H}_{t}-\dot{W}_{t}=-\mathbf{F}\circ\dot{\mathbf{X}}_{t}\,.$ (7)
Converting the Stratonovich differential into Itô [43] and inserting Eq. (2), we find an implicit Langevin equation for the heat in terms of the solutions $\mathbf{X}_{t}$ of Eq. (2),
$\dot{Q}_{t}=-\mathbf{F}\cdot\boldsymbol{\mu}\cdot\mathbf{F}-{\boldsymbol{\nabla}}\cdot(\mathbf{D}\cdot\mathbf{F})-\mathbf{F}\cdot\mathbf{b}\cdot\boldsymbol{\eta}_{t}\,.$ (8)
The above is a natural adaptation of the internal energy, work, and heat to the microscale [31, 32]. A similar extension of the system’s entropy is less trivial. It is common to consider the instantaneous nonequilibrium stochastic entropy [30, 34], $S(\mathbf{x},t)=-\ln p(\mathbf{x},t)$, whose ensemble average is the usual Shannon entropy [48, 49, 50], $\langle S\rangle=-\int{\text{d}}\mathbf{x}\,p(\mathbf{x},t)\ln p(\mathbf{x},t)$. The relation between classical-thermodynamic entropy and stochastic entropy is established for equilibrium states by finding the distribution $p$ that maximizes the Shannon entropy $\langle S\rangle$ [1] (for a given fixed protocol value $\lambda$). Combining the stochastic chain rule [43] and Eqs. (2)–(4), the time-derivative of $S_{t}=S(\mathbf{X}_{t},t)$ is
$\dot{S}_{t}={\boldsymbol{\nabla}}\cdot\left(\frac{\mathbf{J}}{p}\right)-\frac{1}{p}{\boldsymbol{\nabla}}\cdot(\mathbf{D}\cdot{\boldsymbol{\nabla}}p)-[({\boldsymbol{\nabla}}\ln p)\cdot\mathbf{b}]\cdot\boldsymbol{\eta}_{t}\,,$ (9)
where the last two terms are often omitted [51] as they vanish on average.
If a well-defined temperature $T$ exists, it is instructive to consider the thermodynamic heat dissipation $R_{t}$. It quantifies how much extra calorimetric heat $\dot{Q}_{t}$ is wasted in a given irreversible process. According to the second law of thermodynamics, the average thermodynamic dissipation $\langle\dot{R}_{t}\rangle$ is nonnegative [38],
$\dot{R}_{t}\equiv k_{\mathrm{B}}T\dot{S}_{t}-\dot{Q}_{t}\,,\qquad\langle\dot{R}_{t}\rangle\geq 0\,,$ (10)
where $k_{\mathrm{B}}T\dot{S}_{t}$ is the reversible contribution to the heat. Our goal is to arrive at a similar relation in the absence of the conventional temperature $T$.
Informatic entropy production (IEP).—The IEP, $\Sigma_{t}$, quantifies the breaking of time-reversal symmetry in the microstates’ dynamics [30]. We consider a discretized trajectory drawn from Eq. (2), $(\mathbf{X}_{0},\dots,\mathbf{X}_{K\Delta t})=(\mathbf{x}_{0},\dots,\mathbf{x}_{K})$, with an arbitrarily-small timestep $\Delta t$. The probability distribution of obtaining this trajectory under the effect of the protocol sequence $(\lambda_{0},\lambda_{1},\ldots,\lambda_{K})$, $\overrightarrow{\mathcal{P}}\equiv\Pr(\mathbf{x}_{0},\mathbf{x}_{1},\ldots,\mathbf{x}_{K})$, is compared with its reversed counterpart — the probability distribution of obtaining the reverted trajectory, $\overleftarrow{\mathcal{P}}\equiv\Pr(\mathbf{x}_{K},\mathbf{x}_{K-1},\ldots,\mathbf{x}_{0})$, under the reverted protocol $(\lambda_{K},\lambda_{K-1},\ldots,\lambda_{0})$. Explicitly [34, 30],
$\Sigma_{t}=\ln\frac{\overrightarrow{\mathcal{P}}}{\overleftarrow{\mathcal{P}}}\,.$ (11)
For Markovian systems [such as Eq. (2)], it is helpful to decompose the trajectory probabilities using the infinitesimal-time propagator, $\Pr(\mathbf{x}_{k+1},\lambda_{k+1}|\mathbf{x}_{k},\lambda_{k})$, as $\overrightarrow{\mathcal{P}}=p(\mathbf{x}_{0},0)\prod_{k=1}^{K}\Pr(\mathbf{x}_{k},\lambda_{k}|\mathbf{x}_{k-1},\lambda_{k-1})$. This decomposes the IEP rate at time $t=k\Delta t$ into
$\dot{\Sigma}_{t}$ $=$ $\dot{S}_{t}+\dot{\Omega}_{t}\,,$
$\dot{\Omega}_{t}$ $=$ $\lim_{\Delta t\to 0}\frac{1}{\Delta t}\ln\frac{\Pr(\mathbf{x}_{k+1}\,,\lambda_{k+1}|\mathbf{x}_{k},\lambda_{k})}{\Pr(\mathbf{x}_{k},\lambda_{k}|\mathbf{x}_{k+1},\lambda_{k+1})}\,.$ (12)
In Eq. (12) we have identified the time-derivative of the entropy, $\dot{S}_{t}=\ln[p(\mathbf{x}_{k+1},(k+1)\Delta t)/p(\mathbf{x}_{k},k\Delta t)]/\Delta t$, which is related to the Langevin system through Eq. (9). We shall refer to the remainder, $\Omega_{t}$, as the “informatic heat” (IH). The propagator in the denominator, $\Pr(\mathbf{x}_{k},\lambda_{k}|\mathbf{x}_{k+1},\lambda_{k+1})$, is the exact same propagator of Eq. (2), where the jump process is reverted [i.e., the process starts at $\mathbf{x}_{k+1}$ at time $t=k\Delta t$ and ends up at $\mathbf{x}_{k}$ at time $(k+1)\Delta t$ under the protocol value $\lambda_{k+1}$].
Focusing on the IEP has two advantages. First, its ensemble average $\langle\Sigma_{t}\rangle$ is a Kullback-Leibler divergence and thus guaranteed to be nonnegative [50]. Second, the dissipation exactly coincides with the IEP up to $k_{\mathrm{B}}T$, $\dot{R}_{t}=k_{\mathrm{B}}T\dot{\Sigma}_{t}$, in a remarkably wide range of scenarios [30]. In particular, one finds $\dot{Q}_{t}=-k_{\mathrm{B}}T\dot{\Omega}_{t}$ [45] (hence the term “informatic heat” for $\Omega_{t}$). For these cases, therefore, one does not have to resort to simplistic models for the dynamics and then employ Eq. (8) to obtain the thermodynamic heat rate. Instead, one can access the thermodynamic heat rate from the statistics of trajectories [Eqs. (11) and (12)] without any prior knowledge of the underlying equations of motion [52, 53]. Clearly, the correspondence between $\dot{\Omega}_{t},\dot{\Sigma}_{t}$ and $\dot{Q}_{t},R_{t}$ assumes the ER with its well-defined temperature.
More specifically, in Ref. [54] the IH rate is found to satisfy [55]
$\dot{\Omega}_{t}=[(\boldsymbol{\mu}\cdot\mathbf{F})\cdot\mathbf{D}^{-1}]\circ\dot{\mathbf{X}}_{t}.$ (13)
Relating this result to the thermodynamic heat rate, Eq. (7), for any force $\mathbf{F}$, evidently requires the ER. Therefore, if $\mathbf{D}$ and $\boldsymbol{\mu}$ are not related by a scalar, the informatic $\dot{\Sigma}_{t}$ is no longer related to the measurable $\dot{R}_{t}$, even on average. Consequently, the nonnegativity of $\langle\dot{R}_{t}\rangle$ (i.e., the second law) can no longer be established based on trajectory statistics. Below we propose a definition of nonequilibrium temperature that mends this discrepancy.
Nonequilibrium temperature.—We return to the second law, Eq. (10). Noting that it refers only to the first moments of Eqs. (8) and (9), we average over the noise realizations (while using $\langle\boldsymbol{\eta}_{t}\rangle=\mathbf{0}$), and over the instantaneous position, where $\langle[\cdot](\mathbf{X}_{t},\Lambda_{t})\rangle=\int{\text{d}}\mathbf{x}\,p(\mathbf{x},t)[\cdot](\mathbf{x},\Lambda_{t})$. Integration by parts then gives $\langle\dot{Q}_{t}\rangle=-\int{\text{d}}\mathbf{x}\,\mathbf{J}\cdot\mathbf{F}$ and $\langle\dot{S}_{t}\rangle=-\int{\text{d}}\mathbf{x}\,\mathbf{J}\cdot{\boldsymbol{\nabla}}\ln p$. Let us posit that the second law of thermodynamics, Eq. (10), still holds, but with a nonequilibrium temperature $\mathcal{T}$. Using the expressions for $\langle\dot{Q}_{t}\rangle$ and $\langle\dot{S}_{t}\rangle$, we express the average dissipation rate of Eq. (10) as
$\langle\dot{R}_{t}\rangle=k_{\mathrm{B}}\mathcal{T}\int{\text{d}}\mathbf{x}\,\frac{\mathbf{J}}{p}\cdot\mathbf{D}^{-1}\cdot\left(\frac{\mathbf{D}}{k_{\mathrm{B}}\mathcal{T}}\cdot\mathbf{F}p-\mathbf{D}\cdot{\boldsymbol{\nabla}}p\right).$ (14)
This separates out in parentheses a term reminiscent of the probability flux $\mathbf{J}$ of Eq. (4), which depends on our choice of $\mathcal{T}$.
We now impose the following physical conditions on Eq. (14): (i) To satisfy the second law, $\langle\dot{R}_{t}\rangle$ should be a nonnegative functional of $\mathbf{J}$. (ii) To agree with Onsager’s linear-response theory [56, 57], the right-hand side should reduce to a quadratic form of $\mathbf{J}$ near equilibrium and turn the second-law inequality into an equality for reversible processes ($\mathbf{J}\to\mathbf{0}$). (iii) The variable $\mathcal{T}$ should be a configuration-independent scalar. Recalling that the diffusivity is a positive-definite matrix, the simplest choice for the dissipation functional satisfying these three conditions is $\langle\dot{R}_{t}\rangle=k_{\mathrm{B}}\mathcal{T}\int{\text{d}}\mathbf{x}(\mathbf{J}\cdot\mathbf{D}^{-1}\cdot\mathbf{J})/p$ [58]. To complete our construction, we recover the second law of thermodynamics, $\langle\dot{R}_{t}\rangle=k_{\mathrm{B}}\mathcal{T}\langle\dot{S}_{t}\rangle-\langle\dot{Q}_{t}\rangle\geq 0$, if we chose the following drive-dependent nonequilibrium temperature:
$k_{\mathrm{B}}\mathcal{T}=\frac{\int{\text{d}}\mathbf{x}\,\mathbf{J}\cdot\mathbf{F}}{\int{\text{d}}\mathbf{x}\,\mathbf{J}\cdot\mathbf{D}^{-1}\cdot(\boldsymbol{\mu}\cdot\mathbf{F})}\,.$ (15)
Equation (15) is our central result. If the ER holds, such that $\mathbf{D}^{-1}\cdot\boldsymbol{\mu}=(k_{\mathrm{B}}T)^{-1}$, we readily get $\mathcal{T}=T$. Otherwise, the generalized temperature is an intricate quantity. Being drive-dependent, it is evidently less universal than the conventional temperature. This is to be expected for nonequilibrium systems with arbitrarily complex $\mathbf{J}$, $\mathbf{D}$, $\boldsymbol{\mu}$, and $\mathbf{F}$. Furthermore, $\mathcal{T}$ depends on time via the protocol $\Lambda_{t}$ affecting all these quantities. Even for a constant protocol $\Lambda_{t}=\mathrm{const}$, $\mathcal{T}$ will change in time due to the instantaneous distribution $p(\mathbf{x},t)$ entering $\mathbf{J}$. Indeed, nonequilibrium steady states generally have $\mathbf{J}\neq\mathbf{0}$, as nonconservative forces and drift create persistent fluxes across scales. Despite these caveats, Eq. (15) should play an experimentally meaningful role as a generalized temperature in nonequilibrium thermodynamics [19, 17, 29], as discussed below.
Consequences.—First, we reestablish the connection between the IH and heat and between the IEP and dissipation, i.e., between information and thermodynamics. Substituting Eq. (2) in Eq. (13), we find
$\dot{\Omega}_{t}$ $=$ $(\boldsymbol{\mu}\cdot\mathbf{F})\cdot\mathbf{D}^{-1}\cdot(\boldsymbol{\mu}\cdot\mathbf{F})+{\boldsymbol{\nabla}}\cdot(\boldsymbol{\mu}\cdot\mathbf{F})$ (16)
$+[(\boldsymbol{\mu}\cdot\mathbf{F})\cdot\mathbf{D}^{-1}\cdot\mathbf{b}]\cdot\boldsymbol{\eta}_{t}\,,$
which, combined with Eq. (9) and the relation $\dot{\Sigma}_{t}=\dot{S}_{t}+\dot{\Omega}_{t}$, gives the IEP rate as
$\dot{\Sigma}_{t}=\frac{\mathbf{J}}{p}\cdot\mathbf{D}^{-1}\cdot\frac{\mathbf{J}}{p}+\frac{2}{p}{\boldsymbol{\nabla}}\cdot\mathbf{J}+\left(\frac{\mathbf{J}}{p}\cdot\mathbf{D}^{-1}\cdot\mathbf{b}\right)\cdot\boldsymbol{\eta}_{t}\,.$ (17)
Upon taking the averages of Eqs. (16) and (17), we obtain
$k_{\mathrm{B}}\mathcal{T}=-\frac{\langle\dot{Q}_{t}\rangle}{\langle\dot{\Omega}_{t}\rangle}\,,$ (18)
and
$\langle\dot{R}_{t}\rangle=k_{\mathrm{B}}\mathcal{T}\langle\dot{\Sigma}_{t}\rangle=k_{\mathrm{B}}\mathcal{T}\int{\text{d}}\mathbf{x}\frac{\mathbf{J}\cdot\mathbf{D}^{-1}\cdot\mathbf{J}}{p}\,\geq\,0\,.$ (19)
Note that the correspondence holds for the means only and not for the full statistics, $\dot{Q}_{t}\neq-k_{\mathrm{B}}\mathcal{T}\dot{\Omega}_{t}$ and $\dot{R}_{t}\neq k_{\mathrm{B}}\mathcal{T}\dot{\Sigma}_{t}$. For example, the variances are different, $\langle({\text{d}}\Omega_{t})^{2}\rangle_{\mathrm{c}}=2{\text{d}}t\int(\boldsymbol{\mu}\cdot\mathbf{F})\cdot\mathbf{D}\cdot(\boldsymbol{\mu}\cdot\mathbf{F})p{\text{d}}\mathbf{x}$, whereas $\langle({\text{d}}Q_{t})^{2}\rangle_{\mathrm{c}}=2{\text{d}}t\int\mathbf{F}\cdot\mathbf{D}\cdot\mathbf{F}p{\text{d}}\mathbf{x}\neq(k_{\mathrm{B}}\mathcal{T})^{2}\langle({\text{d}}\Omega_{t})^{2}\rangle_{\mathrm{c}}$. (The subscript $\mathrm{c}$ denotes a cumulant.) Since the Jarzynski equality [35, 36] and the Crooks fluctuation theorem [37] require statistics beyond the mean, they are not restored by the generalized $\mathcal{T}$. Thus, without the ER, the various FRs [34] remain related only to informatic quantities (the IEP and IH) [41] and not to thermodynamics (dissipation and heat).
Another relation to be examined is the one between the heat of stochastic systems and the breaking of FDT. Such a connection was established by Harada and Sasa [59], implicitly assuming the ER and a configuration-independent scalar mobility. (We recall that the ER and FDT are equivalent only in equilibrium [60].) Connections found recently for various active systems between FDT breaking and the IH (e.g., Refs. [39, 40]) have similarly made these two strong assumptions. As discussed above, without these assumptions the IH and the thermodynamic heat are separate quantities. In addition, the FDT is violated [61]. It is then unclear which heat is connected to FDT breaking — the IH or the thermodynamic heat? We find that neither connection can be established in the general case [61], implying that the Harada-Sasa relation breaks down as well.
Nevertheless, generalized versions of the FDT and the Harada-Sasa relation can be derived in the absence of the ER. Here we outline these results; see the Supplementary Material (SM) for details [61]. Instead of considering the response to an impulsive linear potential gradient as was done in Ref. [59], we consider a perturbation of the force of the form $\delta\mathbf{F}(\mathbf{x},s)=-\delta h_{\mu}(s)\boldsymbol{\mu}^{-1}(\mathbf{x},s)\cdot\mathbf{D}(\mathbf{x},s)\cdot{\boldsymbol{\nabla}}x_{\mu}$ at time $s<t$, in response to the external perturbation $\delta\mathbf{h}(s)$. The difference between the response of the average force to such perturbation and its correlation with the position are found to be related to the thermodynamic heat [61],
$\langle\dot{Q}_{s}\rangle=\lim_{t\to s^{+}}\left(\frac{{\text{d}}}{{\text{d}}s}\langle\mathbf{F}(\mathbf{X}_{t},t)\cdot\mathbf{X}_{s}\rangle\right.\\ \left.-\frac{\delta}{\delta\mathbf{h}(s)}\cdot\langle\mathbf{F}(\mathbf{X}_{t},t)\rangle\right)\,.$ (20)
If, instead of $\mathbf{F}$, we consider the observable $\mathbf{D}^{-1}\cdot\boldsymbol{\mu}\cdot\mathbf{F}$, we obtain a similar relation for the IH,
$\langle\dot{\Omega}_{s}\rangle=\lim_{t\to s^{+}}\left(\frac{{\text{d}}}{{\text{d}}s}\langle[(\boldsymbol{\mu}\cdot\mathbf{F})\cdot\mathbf{D}](\mathbf{X}_{t},t)\cdot\mathbf{X}_{s}\rangle\right.\\ \left.-\frac{\delta}{\delta\mathbf{h}(s)}\cdot\langle[(\boldsymbol{\mu}\cdot\mathbf{F})\cdot\mathbf{D}](\mathbf{X}_{t},t)\rangle\right)\,.$ (21)
Thus the IH and thermodynamic heat are related to violations of a generalized FDT for different observables. If the ER holds, the two observables coincide and the classical FDT is recovered. With the additional assumption of a configuration-independent scalar mobility the Harada-Sasa relation is recovered as well [61].
We now revisit additional classical-thermodynamic results demonstrating the relevance of the generalized temperature $\mathcal{T}$. More details regarding these consequences are found in the SM [61]. (i) Consider a process that starts and ends at the same protocol value, $\Lambda_{t_{\mathrm{f}}}=\Lambda_{0}$, and then allowed to relax to steady state. Since $\oint_{0}^{\infty}{\text{d}}t\langle\dot{S}_{t}\rangle=0$, we obtain the Clausius inequality, $\oint_{0}^{\infty}{\text{d}}t\langle\dot{Q}_{t}\rangle/(k_{\mathrm{B}}\mathcal{T})=-\oint_{0}^{\infty}{\text{d}}t\langle\dot{\Sigma}_{t}\rangle\leq 0$ [61]. (ii) Construct a Carnot-like cyclic engine consisting of an “isothermal” process (with constant $\mathcal{T}_{\mathrm{H}}$ and $\langle Q_{\mathrm{in}}\rangle>0$), an “adiabatic” process (in which the system “cools” down to $\mathcal{T}_{\mathrm{C}}$), followed by another isothermal process (with constant $\mathcal{T}_{\mathrm{C}}$, and $\langle Q_{\mathrm{out}}\rangle<0$), and an adiabatic process (where the system reaches $\mathcal{T}_{\mathrm{H}}$). The efficiency of such an engine is found as $\langle-W_{\mathrm{cycle}}\rangle/\langle Q_{\mathrm{in}}\rangle=1-\mathcal{T}_{\mathrm{C}}/\mathcal{T}_{\mathrm{H}}+k_{\mathrm{B}}\mathcal{T}_{\mathrm{C}}\langle\Sigma_{\mathrm{cycle}}\rangle/\langle Q_{\mathrm{in}}\rangle$, where $\langle\Sigma_{\mathrm{cycle}}\rangle$ is the mean IEP during a cycle [61]. Thus, the mean IEP quantifies the cycle’s irreversibility compared to a generalized Carnot efficiency, $1-\mathcal{T}_{\mathrm{C}}/\mathcal{T}_{\mathrm{H}}$. (See also Ref. [62] in the context of nonequilibrium engines.) (iii) Define the nonequilibrium free energy, $A_{t}=H_{t}-k_{\mathrm{B}}\mathcal{T}S_{t}$. It is easy to show that it bounds the extractable work in “isothermal” transformations, $\langle-\dot{W}_{t}\rangle=-(\langle\dot{A}_{t}\rangle+k_{\mathrm{B}}\mathcal{T}\langle\dot{\Sigma}_{t}\rangle)\leq-\langle\dot{A}_{t}\rangle$ [61].
An interesting consequence of Eq. (15) relates to a particular class of nonequilibrium systems having multiple “temperatures”, e.g., a mixture of different types of particles, each of which is subjected to a different “thermostat” [63, 64]. These models may be relevant, for instance, for mixtures of passive and active particles [65, 66, 67, 68, 69]. They yield a temperature-like scalar which is a weighted average of the thermostats’ temperatures. Considering a diagonal but nonscalar $\mathbf{D}^{-1}\cdot\boldsymbol{\mu}$, with different inverse temperatures as its eigenvalues, we may view Eq. (15) as a generalization of such a weighted average. Perturbing different global parameters amounts to taking a perturbation $\mathbf{J}$ along different axes, thus picking up different eigenvalues (inverse temperatures) of $\mathbf{D}^{-1}\cdot\boldsymbol{\mu}$. We show a simplistic example in the SM [61].
Conclusion.—We have analyzed a generic nonequilibrium Langevin system [Eq. (2)] which violates the ER [Eq. (1)]. This describes ubiquitous scenarios in driven and active materials since the relation between dissipation (mobility) and fluctuations (diffusivity) is a particular consequence of the equilibrium Boltzmann distribution [15]. We have demonstrated the far-reaching consequences of ER violation, which may be summarized as the detachment of informatic quantities (IH, IEP) from their physical thermodynamic counterparts (heat, dissipation) and the resulting invalidation of fluctuation theorems which rely on these correspondences.
To restore the information-thermodynamics correspondence we have imposed the second law of thermodynamics directly and identified a generalized temperature $\mathcal{T}$ [Eq. (15)], which gives all the thermodynamically required properties: nonnegative dissipation, zero dissipation at equilibrium, linear thermodynamic stability, and the temperature being a macroscopic (integrated) property. The generalized temperature is model-independent — with a calorimetric measurement of the heat and path-sampling for the IH, one may find the system’s $\mathcal{T}$ with no required knowledge of the underlying dynamics. The variety of tools available for the estimation of IEP from trajectories (e.g., sampling marginal distributions [70, 71, 53] or the thermodynamic uncertainty relations [72, 73]) are particularly useful, since in our construction the average IEP remains connected to the average thermodynamic dissipation. The generalized temperature does not restore the various FRs [35, 37, 34, 59], which in the absence of the ER no longer dictate the statistics of thermodynamic properties [41].
Our choice of generalized temperature in Eq. (15) is not unique. We could have chosen $1/(k_{\mathrm{B}}\mathcal{T})=[\langle\dot{S}_{t}\rangle-f(\mathbf{J})]/\langle\dot{Q}_{t}\rangle$ with any function $f(\mathbf{J})$ satisfying $f(\mathbf{J})\sim J^{2}$ for small $J$ and $f(\mathbf{J}\neq\mathbf{0})>0$. All those temperatures will produce the second law and its consequences and coincide with the conventional temperature $T$ when the ER holds. Our definition of $\mathcal{T}$ stems from the simple quadratic form chosen for the dissipation $R_{t}$, Eq. (19). Beside its simplicity, this choice yields the natural and useful correspondence $\langle\dot{Q}_{t}\rangle=-k_{\mathrm{B}}\mathcal{T}\langle\dot{\Omega}_{t}\rangle$.
The temperature $\mathcal{T}$ as presented here is a property of a single closed system exchanging heat with an arbitrary bath. It dictates the flow of heat into or out of the system based on the change in the system’s entropy. Whether $\mathcal{T}$ functions as an effective temperature in other respects remains to be explored. For example, for two systems in contact (i.e., immersed in a bath that drives each of them out of equilibrium), heat should flow from the system with the higher $\mathcal{T}$ to the one with the lower $\mathcal{T}$. These additional attributes are left for a future study.
The applicability of $\mathcal{T}$ as a nonequilibrium temperature can be put to test in simple experimental scenarios. For example, one can make the mobility and diffusivity separately isotropic or anisotropic by controlling, respectively, the particle shape and athermal randomized forces. This will break the ER and allow the tuning of $\mathcal{T}$ according to Eq. (15). One can then study various manifestations of the second law based on $\mathcal{T}$ and its spatial variation [29].
Acknowledgements.
We greatly benefited from discussions with David Dean and Yael Roichman. T.M. acknowledges funding from the Israel Science Foundation (Grant No. 1356/22).

### v2

@@
 Consequently, fluctuation relations for the mechanical work, such as the Jarzynski and Crooks theorems, are invalid.
-We relate the breaking of the correspondence between entropy production and heat dissipation to the breaking of the fluctuation-dissipation theorem.
+We relate the breaking of the correspondence between entropy production and heat dissipation to departure from the fluctuation-dissipation theorem.
 We propose a temperature-like variable which restores this correspondence and gives rise to a generalized second law of thermodynamics, whereby the dissipated heat is necessarily nonnegative and vanishes at equilibrium.
-The Clausius inequality, Carnot maximum efficiency theorem, and relation between the maximum extractable work and the change of free energy are recovered as well.
+The Clausius inequality, Carnot maximum efficiency theorem, and relation between the extractable work and the change of free energy are recovered as well.
 Classical thermodynamics is the central theory for the possible forms of energy transfer in materials.
-It was historically formulated for the purpose of estimating the work that can be extracted from heat engines, utilizing the principle that heat spontaneously flows from “hotter” to “colder” objects.
-This is one version of the second law of thermodynamics, which determines the direction of spontaneous processes.
-Comparing the efficiencies of different heat engines provides a definition of a consistent “hotness” scale which is a thermodynamic state function, i.e., the temperature $T$ [1, 2].
-In light of its vast applicability and usefulness for characterizing the behaviors of equilibrium materials, there is an ongoing effort to extend the theory of thermodynamics to systems that are constantly driven out of equilibrium, for example, self-propelled particles [3, 4, 5], chemical reactions [6, 7], and living systems [8].
+Its many useful consequences, including the Carnot efficiency theorem and the bounds on extractable work via changes in free energy, establish nontrivial and universal relations based on a few macroscopic observables.
+Among these variables is the temperature, which ties microscopic statistics of equilibrium systems to their energetics, e.g., through the equipartition principle or by imposing the mean energy as a constraint for entropy maximization.
+In light of its vast applicability for characterizing the behaviors of equilibrium materials, there is an ongoing effort to extend thermodynamics to systems that are constantly driven out of equilibrium, for example, self-propelled particles [1, 2, 3], sustained chemical reactions [4, 5], and living systems [6].
 Among the complications associated with nonequilibrium systems is the inadequacy of temperature and other variables as thermodynamic state functions [7, 8, 9, 10, 11, 12].
@@
 $\mathbf{D}=k_{\mathrm{B}}T\boldsymbol{\mu}\,.$ (1)
-The diffusivity $\mathbf{D}$ and mobility $\boldsymbol{\mu}$ characterize, respectively, the strength of thermal fluctuations and the dissipative attenuation of motion.
+The diffusivity $\mathbf{D}$ and mobility $\boldsymbol{\mu}$ characterize, respectively, the strength of thermal fluctuations and dissipative attenuation of motion.
 In general they may be space-dependent matrices, reflecting translational and rotational symmetry breaking.
@@
 (1).
-However, in cases where the forces are not conservative [16, 9], or if the fluctuations arise from coarse-graining of fast athermal or active forces [17, 18, 11, 19, 20, 21, 22], there should not be a generic relation between fluctuations (diffusivity) and dissipation (mobility).
-Thus, more often than not, in far-from-equilibrium cases such as active biopolymer networks [23, 24, 25], bacterial swarms [26, 27], and chemical reactions out of equilibrium [28, 7], the ER does not hold.
-In cases where $\mathbf{D}$ and $\boldsymbol{\mu}$ are proportional to one another, one can use Eq.
-(1) to define an effective temperature [2, 11, 17, 29].
-However, in the general case these two position-dependent matrices need not be related by a constant uniform scalar, and Eq.
+However, in cases where forces are not conservative [8, 15], or if fluctuations arise from coarse-graining of fast athermal or active forces [16, 17, 18, 19, 20, 10, 21, 22, 23, 24], there should not be a generic relation between fluctuations (diffusivity) and dissipation (mobility).
+Thus, more often than not, in far-from-equilibrium cases such as active biopolymer networks [25, 26, 27], bacterial swarms [28, 29], and driven chemical reactions [30, 5], the ER does not hold.
+In cases where $\mathbf{D}$ is proportional to $\boldsymbol{\mu}$, one can use Eq.
+(1) to define an effective temperature [7, 10, 31, 19, 18].
+However, in the general case, these two position-dependent tensors are not related by a constant uniform scalar, and Eq.
 (1) is violated.
 One of the theories addressing nonequilibrium systems is stochastic thermodynamics [32].
-By considering Langevin equations or discrete Markov chains it has succeeded in establishing trajectory-wise energy conservation (leading to the first law of thermodynamics) [31, 32, 33] and proving a wealth of fluctuation relations (FRs) [34], such as the Jarzynski equality [35, 36] and Crooks relation [37].
-The latter also yields the second law of thermodynamics for the average entropy changes and heat increments (with some individual trajectories violating it [38]).
-Yet, implicitly, all the aforementioned FRs and extended laws of thermodynamics have relied on the existence of a temperature, defined through the ER, to connect the microscopic statistics to thermodynamic energy, mechanical and chemical work, and calorimentric heat.
-Several definitions of a generalized nonequilibrium temperature were proposed based on generalizations of the fluctuation-dissipation theorem (FDT) [39, 11, 40, 18].
-Most of these definitions still assume a thermal bath that obeys the ER.
-In other cases the ER is required to keep the extended FRs connected to physical measurables.
+By considering Langevin equations or discrete Markov chains it establishes trajectory-wise energy conservation (leading to the first law of thermodynamics) [33, 34, 35] and gives a wealth of fluctuation relations (FRs) [36], such as the Jarzynski and Crooks relations [37, 38, 39].
+The latter also yields the second law of thermodynamics for the average entropy changes and heat increments.
+Yet, implicitly, all the aforementioned FRs and extended laws of thermodynamics have relied on the existence of a temperature, defined through the ER, to connect microscopic statistics to thermodynamic energy, mechanical and chemical work, and calorimetric heat.
+Several definitions of a nonequilibrium temperature were proposed based on generalizations of the fluctuation-dissipation theorem (FDT) and FRs in active matter [10, 20, 40, 41], systems with memory [16, 17], and anomalous diffusion [42, 43].
+Most of these require a thermal bath that obeys the ER.
+In this Letter, we adopt an alternative approach, which is not derived from extentions of FRs or FDT, to identify a temperature-like variable directly, such that the classical-thermodynamic results are recovered.
+In some cases, the ER is required to keep the extended FRs connected to physical measurables.
 For example, the Hatano-Sasa relation [44], an extension of the Jarzynski equality, requires Markovian dynamics only.
@@
 Otherwise, they demand knowledge of the complete microstates’ steady-state distribution, which is usually inaccessible or requires simplistic models [45].
-The generalized temperature proposed here, arising directly from thermodynamics, does not rely on those assumptions.
-We have two aims: (i) to investigate the fate of existing nonequilibrium FRs and extensions to the second law of thermodynamics once the assumption of an ER-based temperature is removed; and (ii) to formulate a generalized, ER-independent temperature and derive the resulting alternative thermodynamic identities.
-Overdamped stochastic energetics.—We consider an ensemble of $N$ interacting degrees of freedom (particles), subjected to a deterministic external protocol $\Lambda_{t}$, which is a prechosen time-varying system parameter.
+The temperature-like variable proposed here, arising directly from thermodynamics, does not rely on those assumptions.
+We have two aims: (i) to investigate the fate of existing nonequilibrium FRs and extensions to the second law of thermodynamics once ER is violated; and (ii) to formulate a generalized, ER-independent temperature and derive the resulting alternative thermodynamic identities.
+Overdamped stochastic energetics.—We consider an ensemble of $N$ interacting degrees of freedom (particles), subjected to a deterministic external time-dependent protocol, $\Lambda_{t}$.
 The time-dependent stochastic microstate is an $N$-component vector, $\mathbf{X}_{t}$.
-We couple the particles to an arbitrary medium, which is assumed to provide them with effective mobility $\boldsymbol{\mu}(\mathbf{x},\lambda)$ and diffusivity $\mathbf{D}(\mathbf{x},\lambda)$ tensors, such that the particles follow the overdamped Langevin equation [Eq.
+We couple the particles to an arbitrary medium determining the mobility $\boldsymbol{\mu}(\mathbf{x},\lambda)$ and diffusivity $\mathbf{D}(\mathbf{x},\lambda)$, such that the particles follow the overdamped Langevin equation [Eq.
 (2) below]; no other property of the surrounding medium is specified.
 The aforementioned $\mathbf{X}_{t}$ and $\Lambda_{t}$ denote a particular time-dependent realizations of $\mathbf{x}$ and $\lambda$.
-The particles are subjected to an additional external force $\mathbf{F}_{\mathrm{ext}}(\mathbf{x},\lambda)$, which may be nonconservative and is not included in the internal energy.
+The particles are subjected to an additional external force $\mathbf{F}_{\mathrm{ext}}(\mathbf{x},\lambda)$, which may be nonconservative and is excluded from the internal energy.
 We denote the total ($N$-component) force acting on the particles by $\mathbf{F}=\mathbf{F}_{\mathrm{ext}}-\boldsymbol{\nabla}H$, where $H(\mathbf{x},\lambda)$ is the system’s Hamiltonian.
-The system is driven out of equilibrium by applying the nonconservative force $\mathbf{F}_{\mathrm{ext}}$ and/or violating the ER [Eq.
+The system is driven out of equilibrium due to the nonconservative force $\mathbf{F}_{\mathrm{ext}}$ and/or violation of ER [Eq.
 (1)].
@@
 where $\mathbf{b}$ is the noise magnitute satisfying $\mathbf{D}=\mathbf{b}\cdot\mathbf{b}^{\mathrm{T}}/2$, $\boldsymbol{\eta}_{t}$ is the normalized thermal noise [$\langle\boldsymbol{\eta}_{t}\rangle=\mathbf{0}$, $\langle\boldsymbol{\eta}_{t}\boldsymbol{\eta}_{t^{\prime}}\rangle=\mathbf{I}\delta(t-t^{\prime})$ where $\mathbf{I}$ is the identity matrix], and the equation is written in the Itô convention [46].
-The so-called “spurious drift” term ${\boldsymbol{\nabla}}\cdot\mathbf{D}$ ensures that the Boltzmann distribution is obtained at equilibrium, i.e., in systems with a well-defined temperature, Eq.
-(1), under the effect of conservative forces [44, 45].
+The “spurious drift” term ${\boldsymbol{\nabla}}\cdot\mathbf{D}$ ensures that Boltzmann distribution is obtained at equilibrium, i.e., in systems with a well-defined temperature, Eq.
+(1), under conservative forces [47, 48].
 The corresponding Fokker-Planck equation [49] for the microstates distribution $p(\mathbf{x},t)$ is
@@
 where $\mathbf{J}$ is the probability flux.
-We now follow Sekimoto’s approach [31, 32] to identify the internal energy, work, and heat for trajectories in generic Langevin systems. The time-derivative of the Hamiltonian is [43]
+We follow Sekimoto’s approach [33, 34] to identify the internal energy, work, and heat for trajectories in generic Langevin systems. The time-derivative of the Hamiltonian is [46]
 $\dot{H}_{t}={\boldsymbol{\nabla}}H\circ\dot{\mathbf{X}}_{t}+\frac{\partial H}{\partial\lambda}\dot{\Lambda}_{t}\,,$ (5)
 where $\circ$ is the Stratonovich product [50].
-The term $(\partial H/\partial\lambda){\text{d}}\Lambda_{t}$ is the external work necessary to change the system’s internal energy over time ${\text{d}}t$.
-The force $\mathbf{F}_{\mathrm{ext}}(\mathbf{X}_{t},\Lambda_{t})$ acting along a displacement ${\text{d}}\mathbf{X}_{t}$ during ${\text{d}}t$ performs an additional work, $\mathbf{F}_{\mathrm{ext}}\circ{\text{d}}\mathbf{X}_{t}$.
-Thus the total work rate is
+$(\partial H/\partial\lambda){\text{d}}\Lambda_{t}$ is the external work necessary to change the system’s internal energy over time ${\text{d}}t$.
+The force $\mathbf{F}_{\mathrm{ext}}(\mathbf{X}_{t},\Lambda_{t})$ acting along a displacement ${\text{d}}\mathbf{X}_{t}$ during ${\text{d}}t$ performs additional work, $\mathbf{F}_{\mathrm{ext}}\circ{\text{d}}\mathbf{X}_{t}$.
+Thus, the total work rate is
 $\dot{W}_{t}=\frac{\partial H}{\partial\lambda}\dot{\Lambda}_{t}+\mathbf{F}_{\mathrm{ext}}\circ\dot{\mathbf{X}}_{t}\,.$ (6)
-According to the first law of thermodynamics, the heat rate at time $t$ is the remaining energy flow into the system,
+According to the first law of thermodynamics, the heat rate is the remaining energy flow into the system,
 $\dot{Q}_{t}=\dot{H}_{t}-\dot{W}_{t}=-\mathbf{F}\circ\dot{\mathbf{X}}_{t}\,.$ (7)
@@
 A similar extension of the system’s entropy is less trivial.
-It is common to consider the instantaneous nonequilibrium stochastic entropy [30, 34], $S(\mathbf{x},t)=-\ln p(\mathbf{x},t)$, whose ensemble average is the usual Shannon entropy [48, 49, 50], $\langle S\rangle=-\int{\text{d}}\mathbf{x}\,p(\mathbf{x},t)\ln p(\mathbf{x},t)$.
-The relation between classical-thermodynamic entropy and stochastic entropy is established for equilibrium states by finding the distribution $p$ that maximizes the Shannon entropy $\langle S\rangle$ [1] (for a given fixed protocol value $\lambda$).
+Consider the instantaneous nonequilibrium stochastic entropy [32, 36], $S(\mathbf{x},t)=-\ln p(\mathbf{x},t)$, whose ensemble average is the Shannon entropy [51, 52, 53], $\langle S\rangle=-\int{\text{d}}\mathbf{x}\,p(\mathbf{x},t)\ln p(\mathbf{x},t)$.
+The relation between classical-thermodynamic entropy and stochastic entropy is established for equilibrium states by finding the distribution $p$ that maximizes $\langle S\rangle$ [54] (for a fixed protocol value $\lambda$).
 Combining the stochastic chain rule [46] and Eqs.
-(2)–(4), the time-derivative of $S_{t}=S(\mathbf{X}_{t},t)$ is
+(2)–(4) yields the time-derivative of $S_{t}=S(\mathbf{X}_{t},t)$,
 $\dot{S}_{t}={\boldsymbol{\nabla}}\cdot\left(\frac{\mathbf{J}}{p}\right)-\frac{1}{p}{\boldsymbol{\nabla}}\cdot(\mathbf{D}\cdot{\boldsymbol{\nabla}}p)-[({\boldsymbol{\nabla}}\ln p)\cdot\mathbf{b}]\cdot\boldsymbol{\eta}_{t}\,,$ (9)
 where the last two terms are often omitted [55] as they vanish on average.
-If a well-defined temperature $T$ exists, it is instructive to consider the thermodynamic heat dissipation $R_{t}$.
-It quantifies how much extra calorimetric heat $\dot{Q}_{t}$ is wasted in a given irreversible process.
+If a well-defined temperature $T$ exists, it is instructive to consider the thermodynamic heat dissipation, $k_{\mathrm{B}}TR_{t}$.
+It quantifies how much extra calorimetric heat $\dot{Q}_{t}$ is wasted in an irreversible process.
 According to the second law of thermodynamics, the average thermodynamic dissipation $\langle\dot{R}_{t}\rangle$ is nonnegative [56],
-$\dot{R}_{t}\equiv k_{\mathrm{B}}T\dot{S}_{t}-\dot{Q}_{t}\,,\qquad\langle\dot{R}_{t}\rangle\geq 0\,,$ (10)
-where $k_{\mathrm{B}}T\dot{S}_{t}$ is the reversible contribution to the heat. Our goal is to arrive at a similar relation in the absence of the conventional temperature $T$.
+$\dot{R}_{t}\equiv\dot{S}_{t}-\frac{\dot{Q}_{t}}{k_{\mathrm{B}}T}\,,\qquad\langle\dot{R}_{t}\rangle\geq 0\,.$ (10)
+Recall that $k_{\mathrm{B}}T\dot{S}_{t}$ is the reversible contribution to the heat. Our goal is to arrive at a similar relation in the absence of the conventional temperature $T$.
 Informatic entropy production (IEP).—The IEP, $\Sigma_{t}$, quantifies the breaking of time-reversal symmetry in the microstates’ dynamics [32].
-We consider a discretized trajectory drawn from Eq.
-(2), $(\mathbf{X}_{0},\dots,\mathbf{X}_{K\Delta t})=(\mathbf{x}_{0},\dots,\mathbf{x}_{K})$, with an arbitrarily-small timestep $\Delta t$.
-The probability distribution of obtaining this trajectory under the effect of the protocol sequence $(\lambda_{0},\lambda_{1},\ldots,\lambda_{K})$, $\overrightarrow{\mathcal{P}}\equiv\Pr(\mathbf{x}_{0},\mathbf{x}_{1},\ldots,\mathbf{x}_{K})$, is compared with its reversed counterpart — the probability distribution of obtaining the reverted trajectory, $\overleftarrow{\mathcal{P}}\equiv\Pr(\mathbf{x}_{K},\mathbf{x}_{K-1},\ldots,\mathbf{x}_{0})$, under the reverted protocol $(\lambda_{K},\lambda_{K-1},\ldots,\lambda_{0})$.
+Consider a discretized trajectory drawn from Eq.
+(2), $(\mathbf{X}_{0},\dots,\mathbf{X}_{K\Delta t})=(\mathbf{x}_{0},\dots,\mathbf{x}_{K})$, with arbitrarily-small timestep $\Delta t$.
+The probability distribution of obtaining this trajectory under the protocol sequence $(\lambda_{0},\lambda_{1},\ldots,\lambda_{K})$, $\overrightarrow{\mathcal{P}}\equiv\Pr(\mathbf{x}_{0},\mathbf{x}_{1},\ldots,\mathbf{x}_{K})$, is compared with the probability distribution of obtaining the reverted trajectory, $\overleftarrow{\mathcal{P}}\equiv\Pr(\mathbf{x}_{K},\mathbf{x}_{K-1},\ldots,\mathbf{x}_{0})$, under the reverted protocol $(\lambda_{K},\lambda_{K-1},\ldots,\lambda_{0})$.
 Explicitly [36, 32],
@@
 In Eq.
-(12) we have identified the time-derivative of the entropy, $\dot{S}_{t}=\ln[p(\mathbf{x}_{k+1},(k+1)\Delta t)/p(\mathbf{x}_{k},k\Delta t)]/\Delta t$, which is related to the Langevin system through Eq.
+(12) we identify the time-derivative of the entropy, $\dot{S}_{t}=\ln[p(\mathbf{x}_{k+1},(k+1)\Delta t)/p(\mathbf{x}_{k},k\Delta t)]/\Delta t$, given by Eq.
 (9).
-We shall refer to the remainder, $\Omega_{t}$, as the “informatic heat” (IH).
+We refer to the remainder, $\Omega_{t}$, as the “informatic heat” (IH).
 The propagator in the denominator, $\Pr(\mathbf{x}_{k},\lambda_{k}|\mathbf{x}_{k+1},\lambda_{k+1})$, is the exact same propagator of Eq.
 (2), where the jump process is reverted [i.e., the process starts at $\mathbf{x}_{k+1}$ at time $t=k\Delta t$ and ends up at $\mathbf{x}_{k}$ at time $(k+1)\Delta t$ under the protocol value $\lambda_{k+1}$].
-Focusing on the IEP has two advantages.
-First, its ensemble average $\langle\Sigma_{t}\rangle$ is a Kullback-Leibler divergence and thus guaranteed to be nonnegative [50].
-Second, the dissipation exactly coincides with the IEP up to $k_{\mathrm{B}}T$, $\dot{R}_{t}=k_{\mathrm{B}}T\dot{\Sigma}_{t}$, in a remarkably wide range of scenarios [30].
-In particular, one finds $\dot{Q}_{t}=-k_{\mathrm{B}}T\dot{\Omega}_{t}$ [45] (hence the term “informatic heat” for $\Omega_{t}$).
-For these cases, therefore, one does not have to resort to simplistic models for the dynamics and then employ Eq.
-(8) to obtain the thermodynamic heat rate.
-Instead, one can access the thermodynamic heat rate from the statistics of trajectories [Eqs.
-(11) and (12)] without any prior knowledge of the underlying equations of motion [52, 53].
-Clearly, the correspondence between $\dot{\Omega}_{t},\dot{\Sigma}_{t}$ and $\dot{Q}_{t},R_{t}$ assumes the ER with its well-defined temperature.
+[…]
 More specifically, in Ref. [59] the IH rate is found to satisfy [60]
 $\dot{\Omega}_{t}=[(\boldsymbol{\mu}\cdot\mathbf{F})\cdot\mathbf{D}^{-1}]\circ\dot{\mathbf{X}}_{t}.$ (13)
-Relating this result to the thermodynamic heat rate, Eq.
-(7), for any force $\mathbf{F}$, evidently requires the ER.
 Therefore, if $\mathbf{D}$ and $\boldsymbol{\mu}$ are not related by a scalar, the informatic $\dot{\Sigma}_{t}$ is no longer related to the measurable $\dot{R}_{t}$, even on average.
 Consequently, the nonnegativity of $\langle\dot{R}_{t}\rangle$ (i.e., the second law) can no longer be established based on trajectory statistics.
-Below we propose a definition of nonequilibrium temperature that mends this discrepancy.
+Below we propose a definition of a nonequilibrium temperature that mends this discrepancy.
 Nonequilibrium temperature.—We return to the second law, Eq.
 (10).
-Noting that it refers only to the first moments of Eqs.
+Noting that it concerns only the means of Eqs.
 (8) and (9), we average over the noise realizations (while using $\langle\boldsymbol{\eta}_{t}\rangle=\mathbf{0}$), and over the instantaneous position, where $\langle[\cdot](\mathbf{X}_{t},\Lambda_{t})\rangle=\int{\text{d}}\mathbf{x}\,p(\mathbf{x},t)[\cdot](\mathbf{x},\Lambda_{t})$.
-Integration by parts then gives $\langle\dot{Q}_{t}\rangle=-\int{\text{d}}\mathbf{x}\,\mathbf{J}\cdot\mathbf{F}$ and $\langle\dot{S}_{t}\rangle=-\int{\text{d}}\mathbf{x}\,\mathbf{J}\cdot{\boldsymbol{\nabla}}\ln p$.
+Integration by parts gives $\langle\dot{Q}_{t}\rangle=-\int{\text{d}}\mathbf{x}\,\mathbf{J}\cdot\mathbf{F}$ and $\langle\dot{S}_{t}\rangle=-\int{\text{d}}\mathbf{x}\,\mathbf{J}\cdot{\boldsymbol{\nabla}}\ln p$.
 Let us posit that the second law of thermodynamics, Eq.
@@
 (10) as
-$\langle\dot{R}_{t}\rangle=k_{\mathrm{B}}\mathcal{T}\int{\text{d}}\mathbf{x}\,\frac{\mathbf{J}}{p}\cdot\mathbf{D}^{-1}\cdot\left(\frac{\mathbf{D}}{k_{\mathrm{B}}\mathcal{T}}\cdot\mathbf{F}p-\mathbf{D}\cdot{\boldsymbol{\nabla}}p\right).$ (14)
-This separates out in parentheses a term reminiscent of the probability flux $\mathbf{J}$ of Eq. (4), which depends on our choice of $\mathcal{T}$.
+$\langle\dot{R}_{t}\rangle=\int{\text{d}}\mathbf{x}\,\frac{\mathbf{J}}{p}\cdot\mathbf{D}^{-1}\cdot\left(\frac{\mathbf{D}}{k_{\mathrm{B}}\mathcal{T}}\cdot\mathbf{F}p-\mathbf{D}\cdot{\boldsymbol{\nabla}}p\right).$ (14)
+The term in the parentheses, which depends on our choice of $\mathcal{T}$, is reminiscent of $\mathbf{J}$ [Eq. (4)].
 We now impose the following physical conditions on Eq.
 (14): (i) To satisfy the second law, $\langle\dot{R}_{t}\rangle$ should be a nonnegative functional of $\mathbf{J}$.
-(ii) To agree with Onsager’s linear-response theory [56, 57], the right-hand side should reduce to a quadratic form of $\mathbf{J}$ near equilibrium and turn the second-law inequality into an equality for reversible processes ($\mathbf{J}\to\mathbf{0}$).
+(ii) To agree with Onsager’s theory [61, 62, 63], the right-hand side should reduce to a quadratic form of $\mathbf{J}$ near equilibrium and turn the second-law inequality into an equality for reversible processes ($\mathbf{J}\to\mathbf{0}$).
 (iii) The variable $\mathcal{T}$ should be a configuration-independent scalar.
-Recalling that the diffusivity is a positive-definite matrix, the simplest choice for the dissipation functional satisfying these three conditions is $\langle\dot{R}_{t}\rangle=k_{\mathrm{B}}\mathcal{T}\int{\text{d}}\mathbf{x}(\mathbf{J}\cdot\mathbf{D}^{-1}\cdot\mathbf{J})/p$ [58].
-To complete our construction, we recover the second law of thermodynamics, $\langle\dot{R}_{t}\rangle=k_{\mathrm{B}}\mathcal{T}\langle\dot{S}_{t}\rangle-\langle\dot{Q}_{t}\rangle\geq 0$, if we chose the following drive-dependent nonequilibrium temperature:
+The simplest choice for the dissipation functional that satisfies these three conditions is $\langle\dot{R}_{t}\rangle=\int{\text{d}}\mathbf{x}(\mathbf{J}\cdot\mathbf{D}^{-1}\cdot\mathbf{J})/p$ [64].
+To complete our construction, we solve the equation $\langle\dot{R}_{t}\rangle=\langle\dot{S}_{t}\rangle-\langle\dot{Q}_{t}\rangle/(k_{\mathrm{B}}\mathcal{T})$ for $k_{\mathrm{B}}\mathcal{T}$, and find the following drive-dependent generalized temperature:
 $k_{\mathrm{B}}\mathcal{T}=\frac{\int{\text{d}}\mathbf{x}\,\mathbf{J}\cdot\mathbf{F}}{\int{\text{d}}\mathbf{x}\,\mathbf{J}\cdot\mathbf{D}^{-1}\cdot(\boldsymbol{\mu}\cdot\mathbf{F})}\,.$ (15)
+We have thus recovered the second law of thermodynamics.
 Equation (15) is our central result.
-If the ER holds, such that $\mathbf{D}^{-1}\cdot\boldsymbol{\mu}=(k_{\mathrm{B}}T)^{-1}$, we readily get $\mathcal{T}=T$.
+If ER holds, such that $\mathbf{D}^{-1}\cdot\boldsymbol{\mu}=(k_{\mathrm{B}}T)^{-1}$, we readily get $\mathcal{T}=T$.
 Otherwise, the generalized temperature is an intricate quantity.
-Being drive-dependent, it is evidently less universal than the conventional temperature.
-This is to be expected for nonequilibrium systems with arbitrarily complex $\mathbf{J}$, $\mathbf{D}$, $\boldsymbol{\mu}$, and $\mathbf{F}$.
+This is expected for nonequilibrium systems with arbitrarily complex $\mathbf{J}$, $\mathbf{D}$, $\boldsymbol{\mu}$, and $\mathbf{F}$.
 Furthermore, $\mathcal{T}$ depends on time via the protocol $\Lambda_{t}$ affecting all these quantities.
@@
 Despite these caveats, Eq.
-(15) should play an experimentally meaningful role as a generalized temperature in nonequilibrium thermodynamics [19, 17, 29], as discussed below.
+(15) should play an experimentally meaningful role as a generalized temperature in nonequilibrium thermodynamics [21, 31, 19, 18], as discussed below.
 Consequences.—First, we reestablish the connection between the IH and heat and between the IEP and dissipation, i.e., between information and thermodynamics. Substituting Eq. (2) in Eq. (13), we find
@@
 Upon taking the averages of Eqs. (16) and (17), we obtain
-$k_{\mathrm{B}}\mathcal{T}=-\frac{\langle\dot{Q}_{t}\rangle}{\langle\dot{\Omega}_{t}\rangle}\,,$ (18)
+$k_{\mathrm{B}}\mathcal{T}=-\langle\dot{Q}_{t}\rangle/\langle\dot{\Omega}_{t}\rangle\,,$ (18)
 and
-$\langle\dot{R}_{t}\rangle=k_{\mathrm{B}}\mathcal{T}\langle\dot{\Sigma}_{t}\rangle=k_{\mathrm{B}}\mathcal{T}\int{\text{d}}\mathbf{x}\frac{\mathbf{J}\cdot\mathbf{D}^{-1}\cdot\mathbf{J}}{p}\,\geq\,0\,.$ (19)
-Note that the correspondence holds for the means only and not for the full statistics, $\dot{Q}_{t}\neq-k_{\mathrm{B}}\mathcal{T}\dot{\Omega}_{t}$ and $\dot{R}_{t}\neq k_{\mathrm{B}}\mathcal{T}\dot{\Sigma}_{t}$.
-For example, the variances are different, $\langle({\text{d}}\Omega_{t})^{2}\rangle_{\mathrm{c}}=2{\text{d}}t\int(\boldsymbol{\mu}\cdot\mathbf{F})\cdot\mathbf{D}\cdot(\boldsymbol{\mu}\cdot\mathbf{F})p{\text{d}}\mathbf{x}$, whereas $\langle({\text{d}}Q_{t})^{2}\rangle_{\mathrm{c}}=2{\text{d}}t\int\mathbf{F}\cdot\mathbf{D}\cdot\mathbf{F}p{\text{d}}\mathbf{x}\neq(k_{\mathrm{B}}\mathcal{T})^{2}\langle({\text{d}}\Omega_{t})^{2}\rangle_{\mathrm{c}}$.
+$\langle\dot{R}_{t}\rangle=\langle\dot{\Sigma}_{t}\rangle=\int{\text{d}}\mathbf{x}\left(\mathbf{J}\cdot\mathbf{D}^{-1}\cdot\mathbf{J}\right)/p\,\geq\,0\,.$ (19)
+Note that the correspondence holds for the means only and not for the full statistics, $\dot{Q}_{t}\neq-k_{\mathrm{B}}\mathcal{T}\dot{\Omega}_{t}$ and $\dot{R}_{t}\neq\dot{\Sigma}_{t}$.
+For example, the variances of the former are different, $\langle({\text{d}}\Omega_{t})^{2}\rangle_{\mathrm{c}}=2{\text{d}}t\int(\boldsymbol{\mu}\cdot\mathbf{F})\cdot\mathbf{D}\cdot(\boldsymbol{\mu}\cdot\mathbf{F})p{\text{d}}\mathbf{x}$, whereas $\langle({\text{d}}Q_{t})^{2}\rangle_{\mathrm{c}}=2{\text{d}}t\int\mathbf{F}\cdot\mathbf{D}\cdot\mathbf{F}p{\text{d}}\mathbf{x}\neq(k_{\mathrm{B}}\mathcal{T})^{2}\langle({\text{d}}\Omega_{t})^{2}\rangle_{\mathrm{c}}$.
 (The subscript ${\mathrm{c}}$ denotes a cumulant.) Since the Jarzynski equality [37, 38] and the Crooks fluctuation theorem [39] require statistics beyond the mean, they are not restored by the generalized $\mathcal{T}$.
 Thus, without the ER, the various FRs [36] remain related only to informatic quantities (the IEP and IH) [44] and not to thermodynamics (dissipation and heat).
-Another relation to be examined is the one between the heat of stochastic systems and the breaking of FDT.
+Another relation to be examined is the one between the heat of stochastic systems and the departure from FDT.
 Such a connection was established by Harada and Sasa [65], implicitly assuming the ER and a configuration-independent scalar mobility.
-(We recall that the ER and FDT are equivalent only in equilibrium [60].) Connections found recently for various active systems between FDT breaking and the IH (e.g., Refs.
+(We recall that the ER and FDT are equivalent only in equilibrium [66].) Connections found recently for various active systems between FDT violation and the IH (e.g., Refs.
 [40, 41]) have similarly made these two strong assumptions.
 As discussed above, without these assumptions the IH and the thermodynamic heat are separate quantities.
-In addition, the FDT is violated [61].
-It is then unclear which heat is connected to FDT breaking — the IH or the thermodynamic heat?
+In addition, the assumptions underlying FDT do not apply [67].
+It is then unclear which heat is connected to FDT violation — the IH or the thermodynamic heat?
 We find that neither connection can be established in the general case [67], implying that the Harada-Sasa relation breaks down as well.
@@
 The difference between the response of the average force to such perturbation and its correlation with the position are found to be related to the thermodynamic heat [67],
-$\langle\dot{Q}_{s}\rangle=\lim_{t\to s^{+}}\left(\frac{{\text{d}}}{{\text{d}}s}\langle\mathbf{F}(\mathbf{X}_{t},t)\cdot\mathbf{X}_{s}\rangle\right.\\ \left.-\frac{\delta}{\delta\mathbf{h}(s)}\cdot\langle\mathbf{F}(\mathbf{X}_{t},t)\rangle\right)\,.$ (20)
+$\langle\dot{Q}_{s}\rangle\!=\!-\!\lim_{t\to s^{+}}\!\left(\frac{{\text{d}}}{{\text{d}}s}\langle\mathbf{F}(\mathbf{X}_{t},t)\cdot\mathbf{X}_{s}\rangle+\frac{\delta}{\delta\mathbf{h}(s)}\cdot\langle\mathbf{F}(\mathbf{X}_{t},t)\rangle\right)\,.$
 If, instead of $\mathbf{F}$, we consider the observable $\mathbf{D}^{-1}\cdot\boldsymbol{\mu}\cdot\mathbf{F}$, we obtain a similar relation for the IH,
-$\langle\dot{\Omega}_{s}\rangle=\lim_{t\to s^{+}}\left(\frac{{\text{d}}}{{\text{d}}s}\langle[(\boldsymbol{\mu}\cdot\mathbf{F})\cdot\mathbf{D}](\mathbf{X}_{t},t)\cdot\mathbf{X}_{s}\rangle\right.\\ \left.-\frac{\delta}{\delta\mathbf{h}(s)}\cdot\langle[(\boldsymbol{\mu}\cdot\mathbf{F})\cdot\mathbf{D}](\mathbf{X}_{t},t)\rangle\right)\,.$ (21)
+$\langle\dot{\Omega}_{s}\rangle=\lim_{t\to s^{+}}\left(\frac{{\text{d}}}{{\text{d}}s}\langle[(\boldsymbol{\mu}\cdot\mathbf{F})\cdot\mathbf{D}](\mathbf{X}_{t},t)\cdot\mathbf{X}_{s}\rangle\right.\\ \left.+\frac{\delta}{\delta\mathbf{h}(s)}\cdot\langle[(\boldsymbol{\mu}\cdot\mathbf{F})\cdot\mathbf{D}](\mathbf{X}_{t},t)\rangle\right)\,.$ (21)
 Thus the IH and thermodynamic heat are related to violations of a generalized FDT for different observables.
@@
 [68] in the context of nonequilibrium engines.) (iii) Define the nonequilibrium free energy, $A_{t}=H_{t}-k_{\mathrm{B}}\mathcal{T}S_{t}$.
-It is easy to show that it bounds the extractable work in “isothermal” transformations, $\langle-\dot{W}_{t}\rangle=-(\langle\dot{A}_{t}\rangle+k_{\mathrm{B}}\mathcal{T}\langle\dot{\Sigma}_{t}\rangle)\leq-\langle\dot{A}_{t}\rangle$ [61].
+It is easy to show that it bounds the extractable work in “isothermal” (constant $\mathcal{T}$) transformations, $\langle-\dot{W}_{t}\rangle=-(\langle\dot{A}_{t}\rangle+k_{\mathrm{B}}\mathcal{T}\langle\dot{\Sigma}_{t}\rangle)\leq-\langle\dot{A}_{t}\rangle$ [67].
 An interesting consequence of Eq.
@@
 (15) as a generalization of such a weighted average.
-Perturbing different global parameters amounts to taking a perturbation $\mathbf{J}$ along different axes, thus picking up different eigenvalues (inverse temperatures) of $\mathbf{D}^{-1}\cdot\boldsymbol{\mu}$.
+These systems may attain $\mathbf{J}=\mathbf{0}$, which results in a ‘$/0$ 0’ ambiguity for $\mathcal{T}$.
+However, perturbing different global parameters amounts to taking a perturbation $\mathbf{J}$ along different axes, thus picking up different eigenvalues (inverse temperatures) of $\mathbf{D}^{-1}\cdot\boldsymbol{\mu}$.
 We show a simplistic example in the SM [67].
@@
 The generalized temperature is model-independent — with a calorimetric measurement of the heat and path-sampling for the IH, one may find the system’s $\mathcal{T}$ with no required knowledge of the underlying dynamics.
-The variety of tools available for the estimation of IEP from trajectories (e.g., sampling marginal distributions [70, 71, 53] or the thermodynamic uncertainty relations [72, 73]) are particularly useful, since in our construction the average IEP remains connected to the average thermodynamic dissipation.
+In particular, the variety of tools available for the estimation of IEP from trajectories (e.g., sampling marginal distributions [76, 77, 58] or the thermodynamic uncertainty relations [78, 79]) will be useful, since in our construction the average IEP remains connected to the average thermodynamic dissipation.
 The generalized temperature does not restore the various FRs [37, 39, 36, 65], which in the absence of the ER no longer dictate the statistics of thermodynamic properties [44].
-Our choice of generalized temperature in Eq.
-(15) is not unique.
-We could have chosen $1/(k_{\mathrm{B}}\mathcal{T})=[\langle\dot{S}_{t}\rangle-f(\mathbf{J})]/\langle\dot{Q}_{t}\rangle$ with any function $f(\mathbf{J})$ satisfying $f(\mathbf{J})\sim J^{2}$ for small $J$ and $f(\mathbf{J}\neq\mathbf{0})>0$.
-All those temperatures will produce the second law and its consequences and coincide with the conventional temperature $T$ when the ER holds.
-Our definition of $\mathcal{T}$ stems from the simple quadratic form chosen for the dissipation $R_{t}$, Eq.
+The three conditions which have been used to find $\mathcal{T}$, Eq.
+(15), do not uniquely determine it.
+Within linear irreversible thermodynamics [62, 61, 63], they do guarantee a single generalized temperature, as they impose the exact leading $\mathcal{O}(J^{2})$ term, $\langle\dot{R}_{t}\rangle=\int{\text{d}}\mathbf{x}(\mathbf{J}\cdot\mathbf{D}^{-1}\cdot\mathbf{J})/p$.
+This leading term is the complete dissipation functional when the Einstein relation holds.
+If one additionally requires this dissipation functional beyond linear theory, $\mathcal{T}$ is then uniquely determined by Eq.
 (15).
-Beside its simplicity, this choice yields the natural and useful correspondence $\langle\dot{Q}_{t}\rangle=-k_{\mathrm{B}}\mathcal{T}\langle\dot{\Omega}_{t}\rangle$.
-The temperature $\mathcal{T}$ as presented here is a property of a single closed system exchanging heat with an arbitrary bath.
-It dictates the flow of heat into or out of the system based on the change in the system’s entropy.
-Whether $\mathcal{T}$ functions as an effective temperature in other respects remains to be explored.
-For example, for two systems in contact (i.e., immersed in a bath that drives each of them out of equilibrium), heat should flow from the system with the higher $\mathcal{T}$ to the one with the lower $\mathcal{T}$.
-These additional attributes are left for a future study.
+The fact that $\mathcal{T}$ is the ratio between $\langle\dot{Q}_{t}\rangle$ (thermodynamics) and $\langle\dot{\Omega}_{t}\rangle$ (information) is then a useful consequence of this choice of dissipation functional.
+The temperature $\mathcal{T}$ obtained here is a property of a single closed system exchanging heat with an arbitrary bath.
+It dictates the direction of heat flow into or out of the system based on the change in the system’s entropy.
+The fact that the entropy production is non-negative, $\langle\dot{R}_{t}\rangle\geq 0$, implies a positive $\mathcal{T}$ if heat is leaving the system and a negative one if heat is entering the system at the nonequilibrium steady state.
+At equilibrium, there is no heat flow.
+It remains to be explored whether $\mathcal{T}$ also dictates the direction of heat flow between two systems in contact.
 The applicability of $\mathcal{T}$ as a nonequilibrium temperature can be put to test in simple experimental scenarios.
@@
 We greatly benefited from discussions with David Dean and Yael Roichman. T.M. acknowledges funding from the Israel Science Foundation (Grant No. 1356/22).
+Supplementary Material: Second Law of Thermodynamics without Einstein Relation
+SI Review: Response and correlation functions
+In this section, we recall the classical fluctuation-dissipation theorem [14] and the Harada-Sasa relation [65].
+We do so in order to highlight the assumptions that underlie both equations and demonstrate the importance of the Einstein relation.
+All results presented in this section are known; we briefly repeat them for the sake of the sections that follow.
+SI.1 The Fokker-Planck equation
+In Eq. (3) of the main text we consider the Fokker-Planck equation,
+$\frac{\partial p(\mathbf{x},t)}{\partial t}=-{\boldsymbol{\nabla}}\cdot[\boldsymbol{\mu}(\mathbf{x},t)\cdot\mathbf{F}(\mathbf{x},t)p(\mathbf{x},t)-\mathbf{D}(\mathbf{x},t)\cdot{\boldsymbol{\nabla}}p(\mathbf{x},t)].$ (S1)
+Numerous authors have discussed why ${\boldsymbol{\nabla}}\cdot[\mathbf{D}(x,t)\cdot{\boldsymbol{\nabla}}]$ is the physically-relevant order of derivatives; see, e.g., Refs.
+[46, 47, 48, 59].
+In short, assuming the Einstein relation is satisfied, $\mathbf{D}(\mathbf{x})=k_{\mathrm{B}}T\boldsymbol{\mu}(\mathbf{x})$, and the forces are conservative, $\mathbf{F}(\mathbf{x})=-{\boldsymbol{\nabla}}H(\mathbf{x})$, this order of derivatives is the one that yields the Boltzmann factor, $p(\mathbf{x},t\to\infty)\sim e^{-H(\mathbf{x})/(k_{\mathrm{B}}T)}$, as the steady-state (equilibrium) distribution of the Fokker-Planck equation.
+The spurious drift term that appears in Eq.
+(2) of the main text was included in order to obtain the corresponding Langevin equation in the Itô representation [47].
+For later purposes, it will be convenient to define the Fokker-Planck operator $\hat{L}$,
+$\frac{\partial p(\mathbf{x},t)}{\partial t}=\hat{L}(\mathbf{x},t)p(\mathbf{x},t),\qquad\hat{L}(\mathbf{x},t)\equiv-{\boldsymbol{\nabla}}\cdot[\boldsymbol{\mu}(\mathbf{x},t)\cdot\mathbf{F}(\mathbf{x},t)-\mathbf{D}(\mathbf{x},t)\cdot{\boldsymbol{\nabla}}].$ (S2)
+Its Hermitian conjugate is
+$\hat{L}^{\dagger}(\mathbf{x},t)\equiv[\boldsymbol{\mu}(\mathbf{x},t)\cdot\mathbf{F}(\mathbf{x},t)+{\boldsymbol{\nabla}}\cdot\mathbf{D}(\mathbf{x},t)]\cdot{\boldsymbol{\nabla}}.$ (S3)
+In Eq. (4) of the main text, we identified the probability flux $\mathbf{J}$,
+$\frac{\partial p(\mathbf{x},t)}{\partial t}=-{\boldsymbol{\nabla}}\cdot\mathbf{J}(\mathbf{x},t),\qquad\mathbf{J}(\mathbf{x},t)\equiv\boldsymbol{\mu}(\mathbf{x},t)\cdot\mathbf{F}(\mathbf{x},t)p(\mathbf{x},t)-\mathbf{D}(\mathbf{x},t)\cdot{\boldsymbol{\nabla}}p(\mathbf{x},t).$ (S4)
+SI.2 The fluctuation-dissipation theorem
+All derivations in this Section are based on Ref. [84]. (See also Refs. [85, 49].) We may formally write the Fokker-Planck propagator as
+$\Pr(\mathbf{x},t|\mathbf{y},s)=\exp\left[\int_{s}^{t}{\text{d}}t^{\prime}\hat{L}(\mathbf{x},t^{\prime})\right]\delta(\mathbf{x}-\mathbf{y}).$ (S5)
+With this, we can compute, e.g., the autocorrelation between two configuration-dependent variables, $A(\mathbf{X}_{t},t)$ and $B(\mathbf{X}_{s},s)$,
+$\langle A(\mathbf{X}_{t},t)B(\mathbf{X}_{s},s)\rangle=\int{\text{d}}\mathbf{y}p(\mathbf{y},s)B(\mathbf{y},s)\exp\left[\int_{s}^{t}{\text{d}}t^{\prime}\hat{L}^{\dagger}(\mathbf{y},t^{\prime})\right]A(\mathbf{y},t).$ (S6)
+Recall the following identity under the variation of the operator $\hat{L}\to\hat{L}+\delta\hat{L}$:
+$\delta\exp\left[\int_{s}^{t}{\text{d}}t^{\prime}\hat{L}(\mathbf{x},t^{\prime})\right]=\int_{s}^{t}{\text{d}}s^{\prime}\exp\left[\int_{s^{\prime}}^{t}{\text{d}}t^{\prime\prime}\hat{L}(\mathbf{x},t^{\prime\prime})\right]\delta\hat{L}(\mathbf{x},s^{\prime})\exp\left[\int_{s}^{s^{\prime}}{\text{d}}t^{\prime}\hat{L}(\mathbf{x},t^{\prime})\right].$ (S7)
+Its Hermitian conjugate reads
+$\delta\exp\left[\int_{s}^{t}{\text{d}}t^{\prime}\hat{L}^{\dagger}(\mathbf{x},t^{\prime})\right]=\int_{s}^{t}{\text{d}}s^{\prime}\exp\left[\int^{s^{\prime}}_{s}{\text{d}}t^{\prime}\hat{L}^{\dagger}(\mathbf{x},t^{\prime})\right]\delta\hat{L}^{\dagger}(\mathbf{x},s^{\prime})\exp\left[\int_{s^{\prime}}^{t}{\text{d}}t^{\prime\prime}\hat{L}^{\dagger}(\mathbf{x},t^{\prime\prime})\right].$ (S8)
+From Eq. (S8), one finds for a perturbation $\delta L(x,s^{\prime})=-\delta s\delta(s^{\prime}-s)L(x,s^{\prime})$
+$\frac{{\text{d}}}{{\text{d}}s}\exp\left[\int_{s}^{t}{\text{d}}t^{\prime}\hat{L}^{\dagger}(\mathbf{x},t^{\prime})\right]=-\hat{L}^{\dagger}(\mathbf{x},s)\exp\left[\int_{s}^{t}{\text{d}}t^{\prime}\hat{L}^{\dagger}(\mathbf{x},t^{\prime})\right].$ (S9)
+With this, we proceed to take the $s$-derivative of Eq. (S6). Using Eq. (S9), integrating by parts, and utilizing Eqs. (S2) – (S4), we find that
+$\frac{{\text{d}}}{{\text{d}}s}\langle A(\mathbf{X}_{t},t)B(\mathbf{X}_{s},s)\rangle$ $=$ $\langle A(\mathbf{X}_{t},t)\left.\frac{\partial B}{\partial t}\right|_{\mathbf{X}_{s},s}\rangle+\int{\text{d}}\mathbf{x}\mathbf{J}(\mathbf{x},s)\cdot[{\boldsymbol{\nabla}}B(\mathbf{x},s)]\exp\left[\int_{s}^{t}{\text{d}}t^{\prime}\hat{L}^{\dagger}(\mathbf{x},t^{\prime})\right]A(\mathbf{x},t)$ (S10)
+$+\int{\text{d}}\mathbf{x}p(\mathbf{x},s)[\mathbf{D}(\mathbf{x},s)\cdot{\boldsymbol{\nabla}}B(\mathbf{x},s)]\cdot{\boldsymbol{\nabla}}\left\{\exp\left[\int_{s}^{t}{\text{d}}t^{\prime}\hat{L}^{\dagger}(\mathbf{x},t^{\prime})\right]A(\mathbf{x},t)\right\}.$
+Now, introduce an instantaneous perturbation to the system’s Hamiltonian, $H(\mathbf{x},s)\to H(\mathbf{x},s)+\delta h_{H}(s){\tilde{H}}(\mathbf{x})$, where $\delta h_{H}(s)$ controls the timing of this impulsive change and ${\tilde{H}}(\mathbf{x})$ is the actual change to the potential.
+As a result, the adjoint Fokker-Planck operator [Eq.
+(S3)] is modified as $\delta\hat{L}^{\dagger}(\mathbf{x},s)=\delta h_{H}(s)[-\boldsymbol{\mu}(\mathbf{x},s)\cdot{\boldsymbol{\nabla}}{\tilde{H}}(\mathbf{x})]\cdot{\boldsymbol{\nabla}}$.
+Thus, using Eq.
+(S8), the response of the mean value of the observable $A(\mathbf{X}_{t},t)$ to the perturbation is
+$\frac{\delta\langle A(\mathbf{X}_{t},t)\rangle}{\delta h_{H}(s)}=-\int{\text{d}}\mathbf{x}p(\mathbf{x},s)[\boldsymbol{\mu}(\mathbf{x},s)\cdot{\boldsymbol{\nabla}}{\tilde{H}}(\mathbf{x})]\cdot{\boldsymbol{\nabla}}\left\{\exp\left[\int_{s}^{t}{\text{d}}t^{\prime}\hat{L}^{\dagger}(\mathbf{x},t^{\prime})\right]A(\mathbf{x},t)\right\}.$ (S11)
+In order to obtain the classical fluctuation-dissipation theorem, we now assume that the system was in equilibrium at the time of the perturbation, $s$.
+“Equilibrium”, in fact, implies two distinct assumptions: (1) that Einstein relation holds, $\mathbf{D}(\mathbf{x},s)=k_{\mathrm{B}}T(s)\boldsymbol{\mu}(\mathbf{x},s)$, where $k_{\mathrm{B}}T(s)$ is a uniform scalar temperature, and (2) that the system was at a flux-free steady-state, $\mathbf{J}(\mathbf{x},s)=\mathbf{0}$.
+Upon inserting these assumptions in Eq.
+(S10) (with $B=\tilde{H}$) and comparing with Eq.
+(S11), we obtain the equilibrium fluctuation-dissipation theorem [14, 84]:
+$\frac{{\text{d}}}{{\text{d}}s}\langle A(\mathbf{X}_{t},t){\tilde{H}}(\mathbf{X}_{s})\rangle=-k_{\mathrm{B}}T\frac{\delta\langle A(\mathbf{X}_{t},t)\rangle}{\delta h_{H}(s)}.$ (S12)
+Important to this work is the case in which Einstein relation is invalid at $s$ (due to, e.g., activity), so there is no uniform scalar $k_{\mathrm{B}}T(s)$ relating $\mathbf{D}(\mathbf{x},s)$ to $\boldsymbol{\mu}(\mathbf{x},s)$.
+This means that the response to a perturbed potential [Eq.
+(S11)] may not be related to the correlation with this potential [Eq.
+(S10)].
+Therefore, the fluctuation-dissipation theorem [Eq.
+(S12)] is not satisfied.
+This is hinted by the fact that the correlation and response functions differ in units by a factor of energy units.
+SI.3 The Harada-Sasa relation
+Particularly useful are correlations among positions, $\langle\mathbf{X}_{t}\mathbf{X}_{s}\rangle$.
+Not only are they (comparably-) easy to extract from experiments and simulations, they are connected to the response to a constant potential gradient (induced, e.g., by gravity), $H(\mathbf{x},s)\to H(\mathbf{x},s)+\delta\mathbf{h}_{H}(s)\cdot\mathbf{x}$ along the direction $\delta\mathbf{h}_{H}/|\delta\mathbf{h}_{H}|$.
+The equilibrium fluctuation-dissipation theorem for correlations among positions reads
+$\frac{{\text{d}}}{{\text{d}}s}\langle X^{\mu}_{t}X^{\nu}_{s}\rangle=-k_{\mathrm{B}}T\frac{\delta\langle X_{t}^{\mu}\rangle}{\delta h_{H}^{\nu}(s)},$ (S13)
+where Greek indices denote different spatial directions.
+Harada and Sasa [65] investigated the breaking of the fluctuation-dissipation theorem in systems that are not in equilibrium, namely, those for which $\mathbf{J}(\mathbf{x},s)\neq\mathbf{0}$.
+What they found is that the heat,
+$\langle\dot{Q}_{t}\rangle=-\int{\text{d}}\mathbf{x}\mathbf{J}(\mathbf{x},t)\cdot\mathbf{F}(\mathbf{x},t),$ (S14)
+which incorporates the nonzero flux $\mathbf{J}(\mathbf{x},s)$ in it, is exactly the difference between the correlation and the response functions,
+$\lim_{t\to s^{+}}\left[\frac{{\text{d}}}{{\text{d}}t}\frac{{\text{d}}}{{\text{d}}s}\langle\mathbf{X}_{t}\cdot\mathbf{X}_{s}\rangle+k_{\mathrm{B}}T\frac{{\text{d}}}{{\text{d}}t}\frac{\delta}{\delta\mathbf{h}_{H}(s)}\cdot\langle\mathbf{X}_{t}\rangle\right]=-\mu(s)\langle\dot{Q}_{s}\rangle.$ (S15)
+This can be seen by the fact that both the heat and the second term in Eq.
+(S10) that we have dropped for $\mathbf{J}(\mathbf{x},s)=\mathbf{0}$ take the form $\int{\text{d}}\mathbf{x}\mathbf{J}(\mathbf{x},s)[\cdot](\mathbf{x},s)$.
+In addition to a valid Einstein relation, they had to assume a uniform scalar mobility, $\boldsymbol{\mu}=\mu(s)\mathbf{I}$.
+The usefulness of the Harada-Sasa relation [65] is twofold.
+First, they found that this extra term is in fact a macroscopic and physically-meaningful observable.
+Second, it given an alternative route to measuring the heat instead of via calorimetry, requiring to observe the breaking of the fluctuation-dissipation theorem.
+This is particularly important in meso- and microscopic systems where the measurement of heat is not trivial.
+We now show how to derive the Harada-Sasa relation and the assumptions it requires.
+First, we assume that the Einstein relation is valid at a time instant $s$, $\mathbf{D}(\mathbf{x},s)=k_{\mathrm{B}}T(s)\boldsymbol{\mu}(\mathbf{x},s)$.
+Upon reintroducing the missing term of Eq.
+(S10) to Eq.
+(S13) and contracting the indices $\mu,\nu$ in the latter, under the response to a linear potential gradient considered above, we have
+$\frac{{\text{d}}}{{\text{d}}s}\langle\mathbf{X}_{t}\cdot\mathbf{X}_{s}\rangle+k_{\mathrm{B}}T\frac{\delta}{\delta\mathbf{h}_{H}(s)}\cdot\langle\mathbf{X}_{t}\rangle=\int{\text{d}}\mathbf{x}\mathbf{J}(\mathbf{x},s)\cdot\exp\left[\int_{s}^{t}{\text{d}}t^{\prime}\hat{L}^{\dagger}(\mathbf{x},t^{\prime})\right]\mathbf{x}.$ (S16)
+Taking a time derivative with respect to $t$ using Eq. (S9) and applying $\hat{L}^{\dagger}$ [Eq. (S3)] on $\mathbf{x}$ gives
+$\frac{{\text{d}}}{{\text{d}}t}\frac{{\text{d}}}{{\text{d}}s}\langle\mathbf{X}_{t}\cdot\mathbf{X}_{s}\rangle+k_{\mathrm{B}}T\frac{{\text{d}}}{{\text{d}}t}\frac{\delta}{\delta\mathbf{h}_{H}(s)}\cdot\langle\mathbf{X}_{t}\rangle=\int{\text{d}}\mathbf{x}\mathbf{J}(\mathbf{x},s)\cdot\exp\left[\int_{s}^{t}{\text{d}}t^{\prime}\hat{L}^{\dagger}(\mathbf{x},t^{\prime})\right]\hat{L}^{\dagger}(\mathbf{x},t)\mathbf{x}$ (S17a)
+$\qquad\qquad=\int{\text{d}}\mathbf{x}\mathbf{J}(\mathbf{x},s)\cdot\exp\left[\int_{s}^{t}{\text{d}}t^{\prime}\hat{L}^{\dagger}(\mathbf{x},t^{\prime})\right][\boldsymbol{\mu}(\mathbf{x},s)\cdot\mathbf{F}(\mathbf{x},s)+{\boldsymbol{\nabla}}\cdot\mathbf{D}(\mathbf{x},s)].$ (S17b)
+We then take the limit where $t\to s^{+}$ and insert the Einstein relation,
+$\lim_{t\to s^{+}}\left[\frac{{\text{d}}}{{\text{d}}t}\frac{{\text{d}}}{{\text{d}}s}\langle\mathbf{X}_{t}\cdot\mathbf{X}_{s}\rangle+k_{\mathrm{B}}T\frac{{\text{d}}}{{\text{d}}t}\frac{\delta}{\delta\mathbf{h}_{H}(s)}\cdot\langle\mathbf{X}_{t}\rangle\right]=\int{\text{d}}\mathbf{x}\mathbf{J}(\mathbf{x},s)\cdot\boldsymbol{\mu}(\mathbf{x},s)\cdot\mathbf{F}(\mathbf{x},s)+\int{\text{d}}\mathbf{x}\mathbf{J}(\mathbf{x},s){\boldsymbol{\nabla}}:\mathbf{D}(\mathbf{x},s).$ (S18)
+The right-hand side is not yet the heat of Eq.
+(S14).
+Finally, to get the Harada-Sasa relation, we must assume space-independent and scalar mobility, $\boldsymbol{\mu}=\mu(s)\mathbf{I}$, in addition to the Einstein relation [meaning that $\mathbf{D}=k_{\mathrm{B}}T(s)\mu(s)\mathbf{I}$ as well], which yields the Harada-Sasa relation, Eq.
+(S15).
+Indeed, the original Harada-Sasa relation [65] and its various generalizations (e.g., Refs.
+[20, 40, 41]) all use constant and scalar mobility (and thus diffusivity, due to the Einstein relation) in their equations of motion.
+Assuming time-translation invariance and a valid Einstein relation, it is possible to reproduce the Harada-Sasa relation by considering the observable $({\text{d}}/{\text{d}}t)\mathbf{A}(\mathbf{X}_{t})=\boldsymbol{\mu}(\mathbf{X}_{t})\dot{\mathbf{X}}_{t}$ instead of $({\text{d}}/{\text{d}}t)\mathbf{A}(\mathbf{X}_{t})=\dot{\mathbf{X}}_{t}$ for cases with a non-diagonal and configuration-dependent mobility [86].
+Much like the observables we propose in Sec.
+SII for the case of violated Einstein relation, this observable may be nontrivial to estimate in experiment.
+To conclude, to get the fluctuation-dissipation theorem, we have assumed that the system was in equilibrium before the perturbation, meaning that the Einstein relations holds ($\mathbf{D}=k_{\mathrm{B}}T\boldsymbol{\mu}$) and that the system is flux-free ($\mathbf{J}=\mathbf{0}$).
+The Harada-Sasa relation relaxed the assumption of a flux-free state.
+But at the same time, one must have both the Einstein relation ($\mathbf{D}=k_{\mathrm{B}}T\boldsymbol{\mu}$) and a space-independent scalar diffusivity [$\boldsymbol{\mu}=\mu(s)\mathbf{I}$].
+These are quite stringent assumptions that may considerably limit the range of applicability of the Harada-Sasa relation out of equilibrium.
+SII Proof of Eqs. (20) and (21) of the main text
+In this section, we aim to generalize the fluctuation-dissipation theorem and Harada-Sasa relation to cases in which Einstein relation is invalid, so there is no uniform scalar $k_{\mathrm{B}}T$ relating $\mathbf{D}$ to $\boldsymbol{\mu}$ (due to different anisotropy properties of these two tensors and/or their different dependence on space; see main text).
+This means that even within a flux-free steady state $\mathbf{J}(\mathbf{x},s)=\mathbf{0}$, the response to a perturbed potential [Eq.
+(S11)] may not be related to the correlation with this potential [Eq.
+(S10)], and clearly the classical fluctuation-dissipation theorem does not hold.
+We show how even without the Einstein relation, an equation reminiscent of the Harada-Sasa relation can be obtained for both the heat [Eq.
+(S14)] and the informatic heat,
+$\langle\dot{\Omega}_{t}\rangle=-\int{\text{d}}\mathbf{x}\mathbf{J}(\mathbf{x},t)\mathbf{D}^{-1}(\mathbf{x},t)\cdot\boldsymbol{\mu}(\mathbf{x},t)\cdot\mathbf{F}(\mathbf{x},t).$ (S19)
+SII.1 Extending the fluctuation-dissipation theorem
+We seek to relate correlations to response functions in the absence of Einstein relation.
+In Sec.
+SI.2 we considered a perturbation to the Hamiltonian, $H(\mathbf{x},s)\to H(\mathbf{x},s)+h_{H}(s){\tilde{H}}(\mathbf{x})$.
+As a result, the forces are modified as $\mathbf{F}(\mathbf{x},s)\to\mathbf{F}(\mathbf{x},s)-\delta h_{H}(s){\boldsymbol{\nabla}}H(\mathbf{x})$, and thus the adjoint Fokker-Planck operator changes by $\delta\hat{L}(\mathbf{x},s)=\delta h_{H}[-\boldsymbol{\mu}(\mathbf{x},s)\cdot{\boldsymbol{\nabla}}{\tilde{H}}(\mathbf{x})]$.
+In the latter expression, the Einstein relation allowed us to replace $\boldsymbol{\mu}(\mathbf{x},s)$ with $\mathbf{D}(\mathbf{x},s)$.
+Since we do not wish to assume the Einstein relation, we consider instead a response to a seemingly peculiar perturbation, $\mathbf{F}(\mathbf{x},s)\to\mathbf{F}(\mathbf{x},s)-h_{\Phi}(s)\boldsymbol{\mu}^{-1}(\mathbf{x},s)\cdot\mathbf{D}(\mathbf{x},s)\cdot{\boldsymbol{\nabla}}{\tilde{\Phi}}(\mathbf{x})$, where ${\tilde{\Phi}}$ is some function.
+With this perturbation, $\mathbf{D}$ appears instead of $\boldsymbol{\mu}$ in the variation of the adjoint Fokker-Planck operator, $\delta\hat{L}(\mathbf{x},s)=\delta h_{\Phi}(s)[-\mathbf{D}(\mathbf{x},s)\cdot{\boldsymbol{\nabla}}{\tilde{\Phi}}(\mathbf{x})]$.
+The response function for such perturbation is
+$\frac{\delta\langle A(\mathbf{X}_{t},t)\rangle}{\delta h_{\Phi}(s)}=-\int{\text{d}}\mathbf{x}p(\mathbf{x},s)[\mathbf{D}(\mathbf{x},s)\cdot{\boldsymbol{\nabla}}{\tilde{\Phi}}(\mathbf{x})]\cdot{\boldsymbol{\nabla}}\exp\left[\int_{s}^{t}{\text{d}}t^{\prime}\hat{L}^{\dagger}(\mathbf{x},t^{\prime})\right]A(\mathbf{x},t).$ (S20)
+(The question whether such a perturbation is experimentally realizable is a nontrivial one.)
+In order to obtain a generalized fluctuation-dissipation relation, we now assume that $\mathbf{J}(\mathbf{x},s)=\mathbf{0}$.
+(when this is not the case, we get an extension of the Harada-Sasa relations; see Sec.
+SII.2.) This is the assumption that makes the second term in Eq.
+(S10) disappear as in Sec.
+SI.2.
+In the absence of Einstein relation, these $\mathbf{J}=\mathbf{0}$ steady states are not equilibrium states, and they are the ones that lead to the ambiguous generalized temperature, Eq.
+(15) of the main text; see discussion in the last paragraph before the conclusion section in the main text.
+We give a simplistic example of such a state in Sec.
+SIV below.
+The flux-free steady-state constraint, $\mathbf{J}=\mathbf{0}$, implies $\mathbf{F}=-\boldsymbol{\mu}^{-1}\cdot\mathbf{D}\cdot{\boldsymbol{\nabla}}\Phi$, where $\Phi$ is a potential-like function with which $e^{-\Phi}$ is the steady-state distribution.
+With this in mind, we see that the perturbation proposed above is equivalent to the perturbation $\Phi(\mathbf{x},s)\to\Phi(\mathbf{x},s)+h_{\Phi}(s){\tilde{\Phi}}(\mathbf{x})$.
+Thus, assuming that one begins from a flux-free steady-state [$\mathbf{J}(\mathbf{x},s)=\mathbf{0}$] but without the Einstein relation, upon comparing Eq.
+(S20) to Eq.
+(S10), and choosing $B(\mathbf{X}_{s},s)={\tilde{\Phi}}(\mathbf{X}_{s})$, we find
+$\frac{{\text{d}}}{{\text{d}}s}\langle A(\mathbf{X}_{t},t){\tilde{\Phi}}(\mathbf{X}_{s})\rangle=-\frac{\delta\langle A(\mathbf{X}_{t},t)\rangle}{\delta h_{\Phi}(s)}.$ (S21)
+This expression first appeared in Ref.
+[87].
+As mentioned above, creating a perturbation to $\Phi$ may not be as experimentally realizable as a direct perturbation of the Hamiltonian.
+However, if one could apply a force $\mathbf{F}$ for which the quantity $\mathbf{D}^{-1}(\mathbf{x},s)\cdot\boldsymbol{\mu}(\mathbf{x},s)\cdot\mathbf{F}(\mathbf{x},s)$ (without Einstein relation) is derived from a potential $\Phi$, one could possibly perturb that potential.
+Obviously, if the Einstein relation is obeyed, $\mathbf{D}(\mathbf{x},s)=k_{\mathrm{B}}T(s)\boldsymbol{\mu}(\mathbf{x},s)$, we get back the classical fluctuation-dissipation theorem, Eq.
+(S12).
+SII.2 Extending the Harada-Sasa relation
+We continue by extending the Harada-Sasa relation.
+To do so, we relax the assumption from Sec.
+SII.1 for which the system is initially in a flux-free steady-state, such that now $\mathbf{J}(\mathbf{x},s)\neq\mathbf{0}$.
+As in Sec.
+SII.1, we still choose $B(\mathbf{X}_{s},s)={\tilde{\Phi}}(\mathbf{X}_{s})$ in Eq.
+(S10), which now gives
+$\frac{{\text{d}}}{{\text{d}}s}\langle A(\mathbf{X}_{t},t){\tilde{\Phi}}(\mathbf{X}_{s})\rangle+\frac{\delta\langle A(\mathbf{X}_{t},t)\rangle}{\delta h_{\Phi}(s)}=\int{\text{d}}\mathbf{x}\mathbf{J}(\mathbf{x},s)\cdot[{\boldsymbol{\nabla}}{\tilde{\Phi}}(\mathbf{x},s)]\exp\left[\int_{s}^{t}{\text{d}}t^{\prime}\hat{L}^{\dagger}(\mathbf{x},t^{\prime})\right]A(\mathbf{x},t).$ (S22)
+As we have seen in Eq.
+(S18), we cannot obtain the force through a derivative with respect to $t$.
+Thus, instead of choosing $\mathbf{A}(\mathbf{x},t)=\mathbf{x}$, we choose the force as our observable, $\mathbf{A}(\mathbf{x},t)=\mathbf{F}(\mathbf{x},t)$.
+At the same time, we choose a linear ${\tilde{\Phi}}(\mathbf{x})=\mathbf{x}$, and take the limit where $t$ is a time instant shortly after $s$,
+$\lim_{t\to s^{+}}\left[\frac{{\text{d}}}{{\text{d}}s}\langle\mathbf{X}_{s}\cdot\mathbf{F}(\mathbf{X}_{t},t)\rangle+\frac{\delta}{\delta\mathbf{h}_{\Phi}(s)}\cdot\langle\mathbf{F}(\mathbf{X}_{t},t)\rangle\right]=-\langle\dot{Q}_{s}\rangle.$ (S23)
+This is Eq.
+(20) of the main text.
+Thus, without the Einstein relation, also the heat is related to a peculiar response in addition to the cumbersome correlation.
+On the other hand, by choosing an even less trivial observable, $\mathbf{A}(\mathbf{x},s)=\mathbf{D}^{-1}(\mathbf{x},t)\cdot\boldsymbol{\mu}(\mathbf{x},t)\cdot\mathbf{F}(\mathbf{x},t)$, we find a similar relation for the informatic heat,
+$\lim_{t\to s^{+}}\left[\frac{{\text{d}}}{{\text{d}}s}\langle\mathbf{X}_{s}\cdot\mathbf{D}^{-1}(\mathbf{X}_{t},t)\cdot\boldsymbol{\mu}(\mathbf{X}_{t},t)\cdot\mathbf{F}(\mathbf{X}_{t},t)\rangle+\frac{\delta}{\delta\mathbf{h}_{\Phi}(s)}\cdot\langle\mathbf{D}^{-1}(\mathbf{X}_{t},t)\cdot\boldsymbol{\mu}(\mathbf{X}_{t},t)\cdot\mathbf{F}(\mathbf{X}_{t},t)\rangle\right]=\langle\dot{\Omega}_{s}\rangle.$ (S24)
+This is Eq. (21) of the main text.
+With the Einstein relation, the response is to a conservative force as before, and there is no distinction between the heat $\dot{Q}_{s}$ and the informatic heat $\dot{\Omega}_{s}$ up to the temperature.
+In order to get the usual Harada-Sasa relation [65], one should assume in addition that the mobility is uniform and scalar, $\boldsymbol{\mu}=\mu(s)\mathbf{I}$.
+This can be seen by multiplying both sides of Eq.
+(S23) by $\mu$ [or, conversely, Eq.
+(S24) by $k_{\mathrm{B}}T\mu$] and insert $\mu$ into the ensemble averages on the left-hand side.
+Identifying that $\mu\mathbf{F}(\mathbf{X}_{t},t)$ (note that ${\boldsymbol{\nabla}}\cdot\mathbf{D}$ with $\mathbf{D}=k_{\mathrm{B}}T\mu\mathbf{I}$ is zero) is the average of the velocity $({\text{d}}/{\text{d}}t)\mathbf{X}_{t}$, we converge back to the usual Harada-Sasa relation, Eq.
+(S15).
+SIII Classical-thermodynamic relations
+In this section, we provide further details on the proof and the assumptions of the thermodynamic theorems that result from the extended second law of thermodynamics constructed in the main text,
+$\langle\dot{R}_{t}\rangle=k_{\mathrm{B}}\mathcal{T}\langle\dot{S}_{t}\rangle-\langle\dot{Q}_{t}\rangle\geq 0.$ (S25)
+Reintroducing the protocol $\Lambda_{t}$, we recall the expressions described in the main text for the average heat, change in entropy, and heat dissipation:
+$\langle\dot{Q}_{t}\rangle$ $=$ $-\int{\text{d}}\mathbf{x}\mathbf{J}(\mathbf{x},t)\cdot\mathbf{F}(\mathbf{x},\Lambda_{t})$ (S26)
+$=$ $-\int{\text{d}}\mathbf{x}p(\mathbf{x},t)\left\{\mathbf{F}(\mathbf{x},\Lambda_{t})\cdot\boldsymbol{\mu}(\mathbf{x},\Lambda_{t})\cdot\mathbf{F}(\mathbf{x},\Lambda_{t})+{\boldsymbol{\nabla}}\cdot\left[\mathbf{D}(\mathbf{x},\Lambda_{t})\cdot\mathbf{F}(\mathbf{x},\Lambda_{t})\right]\right\},$
+$\langle\dot{S}_{t}\rangle$ $=$ $-\int{\text{d}}\mathbf{x}\mathbf{J}(\mathbf{x},t)\cdot{\boldsymbol{\nabla}}\ln p(\mathbf{x},t)=\int{\text{d}}\mathbf{x}[{\boldsymbol{\nabla}}\cdot\mathbf{J}(\mathbf{x},t)]\ln p(\mathbf{x},t)$ (S27)
+$=$ $-\int{\text{d}}\mathbf{x}p(\mathbf{x},t)\left\{\mathbf{F}(\mathbf{x},\Lambda_{t})\cdot\boldsymbol{\mu}(\mathbf{x},\Lambda_{t})\cdot\frac{{\boldsymbol{\nabla}}p(\mathbf{x},t)}{p(\mathbf{x},t)}+{\boldsymbol{\nabla}}\cdot\left[\mathbf{D}(\mathbf{x},\Lambda_{t})\cdot\frac{{\boldsymbol{\nabla}}p(\mathbf{x},t)}{p(\mathbf{x},t)}\right]\right\},$
+$\langle\dot{R}_{t}\rangle$ $=$ $k_{\mathrm{B}}\mathcal{T}\int{\text{d}}\mathbf{x}\frac{\mathbf{J}(\mathbf{x},t)\cdot\mathbf{D}^{-1}(\mathbf{x},\Lambda_{t})\cdot\mathbf{J}(\mathbf{x},t)}{p(\mathbf{x},t)}.$ (S28)
+The nonequilibrium temperature that we identified in the main text is
+$\frac{1}{k_{\mathrm{B}}\mathcal{T}(t)}$ $=$ $\frac{\int{\text{d}}\mathbf{x}\mathbf{J}(\mathbf{x},t)\cdot\mathbf{D}^{-1}(\mathbf{x},\Lambda_{t})\cdot[\boldsymbol{\mu}(\mathbf{x},\Lambda_{t})\cdot\mathbf{F}(\mathbf{x},\Lambda_{t})]}{\int{\text{d}}\mathbf{x}\mathbf{J}(\mathbf{x},t)\cdot\mathbf{F}(\mathbf{x},\Lambda_{t})}$ (S29)
+$=$ $\frac{\int{\text{d}}\mathbf{x}p(\mathbf{x},t)\left\{[\boldsymbol{\mu}(\mathbf{x},\Lambda_{t})\cdot\mathbf{F}(\mathbf{x},\Lambda_{t})]\cdot\mathbf{D}^{-1}(\mathbf{x},\Lambda_{t})\cdot[\boldsymbol{\mu}(\mathbf{x},\Lambda_{t})\cdot\mathbf{F}(\mathbf{x},\Lambda_{t})]+{\boldsymbol{\nabla}}\cdot\left[\boldsymbol{\mu}(\mathbf{x},\Lambda_{t})\cdot\mathbf{F}(\mathbf{x},\Lambda_{t})\right]\right\}}{\int{\text{d}}\mathbf{x}p(\mathbf{x},t)\left\{\mathbf{F}(\mathbf{x},\Lambda_{t})\cdot\boldsymbol{\mu}(\mathbf{x},\Lambda_{t})\cdot\mathbf{F}(\mathbf{x},\Lambda_{t})+{\boldsymbol{\nabla}}\cdot\left[\mathbf{D}(\mathbf{x},\Lambda_{t})\cdot\mathbf{F}(\mathbf{x},\Lambda_{t})\right]\right\}}.$
+It depends on both the protocol (through the variation of the forcing $\mathbf{F}$, mobility $\boldsymbol{\mu}$, and diffusivity $\mathbf{D}$) and on time (through the dynamics that the distribution $p$ undergoes given the protocol evolution).
+We identify the connection between informatic entropy production and heat dissipation, $\langle\dot{\Sigma}_{t}\rangle=\langle\dot{R}_{t}\rangle/(k_{\mathrm{B}}\mathcal{T})$, and between the informatic and calorimetric heats, $\langle\dot{\Omega}_{t}\rangle=\langle\dot{Q}_{t}\rangle/(k_{\mathrm{B}}\mathcal{T})$.
+SIII.1 Clausius inequality
+Within classical thermodynamics, the Clausius inequality states that for any transformation (reversible or not) starting and ending at the same equilibrium state, $\oint{\text{d}}Q/T_{\mathrm{env}}\leq 0$, where ${\text{d}}Q$ are the heat increments, $T_{\mathrm{env}}$ is the temperature of the environment, and the closed integral is parameterized with the protocol value.
+We now investigate the fate of this theorem with our generalized temperature.
+Consider a process under a circular protocol $\Lambda_{t}=\lambda^{\circ}(t)$, starting at steady state.
+Quantities with the circle superscript denote that they are computed for this circular protocol.
+Between times $t\in[0,t^{\circ}]$, the protocol changes such that $\lambda^{\circ}(0)=\lambda^{\circ}(t^{\circ})\equiv\lambda^{\circ}_{0}$.
+Once it reaches that value, it remains constant, $\lambda^{\circ}(t>t^{\circ})=\lambda^{\circ}_{0}$, so the system may now decay under the constant protocol value to its corresponding steady-state distribution.
+We assume that the system has a finite relaxation time for that protocol value, which we denote $\tau(\lambda^{\circ}_{0})$.
+Thus, when the protocol reaches $\lambda^{\circ}_{0}$ at time $t^{\circ}$, the system reaches the steady-state distribution (the same as the initial one) only after an additional transient, at time $t=t_{\mathrm{f}}^{\circ}\gg t^{\circ}+\tau(\lambda^{\circ}_{0})$, meaning that $p(\mathbf{x},t_{\mathrm{f}}^{\circ})\to p(\mathbf{x},0)$.
+Since the system begins at and approaches to the same state, the entropy — a state function — does not change during this process,
+$\langle S^{\circ}_{t_{\mathrm{f}}^{\circ}}\rangle-\langle S^{\circ}_{0}\rangle=\int_{0}^{t_{\mathrm{f}}^{\circ}}{\text{d}}t\langle\dot{S}^{\circ}_{t}\rangle=0.$ (S30)
+At every instant during this process, the generalized temperature $\mathcal{T}^{\circ}(t)$ is given by Eq.
+(S29), the average heat rate $\langle\dot{Q}_{t}^{\circ}\rangle$ by Eq.
+(S26), and the average heat dissipation rate $\langle\dot{R}_{t}^{\circ}\rangle=k_{\mathrm{B}}\mathcal{T}^{\circ}(t)\langle\dot{\Sigma}_{t}^{\circ}\rangle$ by Eq.
+(S28).
+Combining that with Eq.
+(S25), we find the nonequilibrium analogue to the Clausius inequality,
+$\int_{0}^{t_{\mathrm{f}}^{\circ}}{\text{d}}t\frac{\langle\dot{Q}^{\circ}_{t}\rangle}{k_{\mathrm{B}}\mathcal{T}^{\circ}(t)}=-\int_{0}^{t_{\mathrm{f}}^{\circ}}{\text{d}}t\frac{\langle\dot{R}^{\circ}_{t}\rangle}{k_{\mathrm{B}}\mathcal{T}^{\circ}(t)}=-\int_{0}^{t_{\mathrm{f}}^{\circ}}{\text{d}}t\langle\dot{\Sigma}^{\circ}_{t}\rangle\leq 0.$ (S31)
+[The integrals in Eqs. (S30) and (S31) are between 00 and $t_{\mathrm{f}}^{\circ}$ because only at $t_{\mathrm{f}}^{\circ}\gg t^{\circ}+\tau(\lambda^{\circ}_{0})$ the system reaches a steady-state.]
+Comparing Eq. (S31) and the classical Clausius inequality, we make a few remarks:
+  • The temperature in the Clausius inequality is of the heat reservoir (which is of infinite size and in equilibrium).
+In our case, other than setting $\mathbf{D}(\mathbf{x},\Lambda_{t})$ and $\boldsymbol{\mu}(\mathbf{x},\Lambda_{t})$, the medium in which the system is immersed does not have any specified properties.
+Therefore, we perceive the nonequilibrium temperature $\mathcal{T}$ as the “system’s” temperature.
+At best, inspired by the zeroth law, we may set the medium’s temperature, for constant $\Lambda_{t}=\lambda$, as the limiting temperature of the system in contact, $\mathcal{T}(t\gg\tau(\lambda))$.
+  • For reversible (near-equilibrium) processes, the Clausius inequality is an equality.
+Here, even if the Einstein relation does not apply, we find that at flux-free steady state, $\mathbf{J}=\mathbf{0}$, the generalized Clausius inequality becomes an equality.
+Moreover, it indicates that the flux-free state is stable (that is, for a small departure from equilibrium $\delta\mathbf{J}$, the dissipation is quadratic in it and hence much smaller).
+Lastly, based on various coarse graining techniques of the informatic entropy production (e.g., the thermodynamic uncertainty relations [78, 79]), if a positive estimate of $\langle\Sigma_{t}\rangle$ is found,one obtains a tighter bound than the usual Clausius inequality.
+SIII.2 Efficiency of a heat engine
+In this section, we consider the extension of the Carnot efficiency theorem with our generalized temperature.
+To this end, we consider $\Lambda_{t}$ that is a combination of at least two protocols.
+Notice that the nonequilibrium temperature $\mathcal{T}(t)$ and average heat rate $\langle\dot{Q}_{t}\rangle$ are determined solely by the imposed protocol $\Lambda_{t}$ and the initial conditions $p(\mathbf{x},0)$.
+This can be seen by inserting $p(\mathbf{x},t)=\exp[\int_{0}^{t}\hat{L}(\mathbf{x},\Lambda_{t})]p(\mathbf{x},0)$ in Eqs.
+(S29) and (S26).
+For a set of initial condition $p(\mathbf{x},0)$, we realize that the condition $\langle\dot{Q}_{t}\rangle=0$ defines an adiabatic manifold in protocol space (e.g., if the forcing is such that $\mathbf{F}(\mathbf{x},\Lambda_{t})\cdot\boldsymbol{\mu}(\mathbf{x},\Lambda_{t})\cdot\mathbf{F}(\mathbf{x},\Lambda_{t})=-{\boldsymbol{\nabla}}\cdot\left[\mathbf{D}(\mathbf{x},\Lambda_{t})\cdot\mathbf{F}(\mathbf{x},\Lambda_{t})\right]$, or if the system’s container is isolated), and the condition $\partial\mathcal{T}(t)/\partial t=0$ gives a first-order differential equation for $\Lambda_{t}$ (under the initial condition of the desired temperature $\mathcal{T}$ for $\Lambda_{0}$) which determines an “isothermal” surface.
+We comment that, under the nonequilibrium setup of the main text, it may be nontrivial to realize such “isothermal” process.
+This is why extending the zeroth law of thermodynamics could be quite beneficial, as then an appropriate (nonequilibrium) environment would enforce a constant temperature, just as having an isolated container ensures an adiabatic process.
+We construct the following cycle out of the initial condition $p(\mathbf{x},0)$:
+  1. 1.
+    An isothermal protocol $\Lambda_{t}=\lambda^{\mathcal{T}}_{\mathrm{H}}(t)$ acting during $t\in[0,t^{\mathcal{T}}_{\mathrm{H}}]$ for which $\mathcal{T}(t)=\mathcal{T}_{\mathrm{H}}=\mathrm{const}$.
+The average accumulated heat in this step is
+    $\langle Q_{\mathrm{in}}\rangle\equiv\int_{0}^{t^{\mathcal{T}}_{\mathrm{H}}}{\text{d}}t\langle\dot{Q}_{t}\rangle.$ (S32)
+    The protocol $\lambda^{\mathcal{T}}_{\mathrm{H}}(t)$ is such that $\langle Q_{\mathrm{in}}\rangle>0$.
+  2. 2.
+    An adiabatic transformation $\Lambda_{t}=\lambda^{Q}_{\mathrm{H}\to\mathrm{C}}(t)$ during $t\in[t^{\mathcal{T}}_{\mathrm{H}},t^{\mathcal{T}}_{\mathrm{H}}+t^{Q}_{\mathrm{H}\to\mathrm{C}}]$ that reaches a state $p(\mathbf{x},t^{\mathcal{T}}_{\mathrm{H}}+t^{Q}_{\mathrm{H}\to\mathrm{C}})$ for which we have $\mathcal{T}(t^{\mathcal{T}}_{\mathrm{H}}+t^{Q}_{\mathrm{H}\to\mathrm{C}})=\mathcal{T}_{\mathrm{C}}$.
+As an adiabatic process, $\langle Q_{t}\rangle=0$ during this step.
+  3. 3.
+    Another isothermal protocol $\Lambda_{t}=\lambda^{\mathcal{T}}_{\mathrm{C}}(t)$ acting during $t\in[t^{\mathcal{T}}_{\mathrm{H}}+t^{Q}_{\mathrm{H}\to\mathrm{C}},t^{\mathcal{T}}_{\mathrm{H}}+t^{Q}_{\mathrm{H}\to\mathrm{C}}+t^{\mathcal{T}}_{\mathrm{C}}]$ for which $\mathcal{T}(t)=\mathcal{T}_{\mathrm{C}}=\mathrm{const}$.
+The average accumulated heat in this step is
+    $\langle Q_{\mathrm{out}}\rangle\equiv\int_{t^{\mathcal{T}}_{\mathrm{H}}+t^{Q}_{\mathrm{H}\to\mathrm{C}}}^{t^{\mathcal{T}}_{\mathrm{H}}+t^{Q}_{\mathrm{H}\to\mathrm{C}}+t^{\mathcal{T}}_{\mathrm{C}}}{\text{d}}t\langle\dot{Q}_{t}\rangle,$ (S33)
+    In order for the engine to be at all useful, the protocol should be tailored such that $\langle Q_{\mathrm{out}}\rangle<0$.
+  4. 4.
+    Another adiabatic transformation $\Lambda_{t}=\lambda^{Q}_{\mathrm{C}\to\mathrm{H}}(t)$ during $t\in[t^{\mathcal{T}}_{\mathrm{H}}+t^{Q}_{\mathrm{H}\to\mathrm{C}}+t^{\mathcal{T}}_{\mathrm{C}},t_{\mathrm{f}}]$, where $t_{\mathrm{f}}=t^{\mathcal{T}}_{\mathrm{H}}+t^{Q}_{\mathrm{H}\to\mathrm{C}}+t^{\mathcal{T}}_{\mathrm{C}}+t^{Q}_{\mathrm{C}\to\mathrm{H}}$, that reaches exactly the initial $p(\mathbf{x},t_{\mathrm{f}})=p(\mathbf{x},0)$ with $\lambda^{Q}_{\mathrm{C}\to\mathrm{H}}(t_{\mathrm{f}})=\lambda^{\mathcal{T}}_{\mathrm{H}}(0)$, such that indeed $\mathcal{T}(t_{\mathrm{f}})=\mathcal{T}_{\mathrm{H}}$.
+As an adiabatic process, $\langle Q_{t}\rangle=0$ during this step as well.
+The total informatic entropy production over the four steps is denoted as
+$\langle\Sigma_{\mathrm{cycle}}\rangle\equiv\int_{0}^{t^{\mathcal{T}}_{\mathrm{H}}}{\text{d}}t\langle\dot{\Sigma}_{t}\rangle+\int_{t^{\mathcal{T}}_{\mathrm{H}}}^{t^{\mathcal{T}}_{\mathrm{H}}+t^{Q}_{\mathrm{H}\to\mathrm{C}}}{\text{d}}t\langle\dot{\Sigma}_{t}\rangle+\int_{t^{\mathcal{T}}_{\mathrm{H}}+t^{Q}_{\mathrm{H}\to\mathrm{C}}}^{t^{\mathcal{T}}_{\mathrm{H}}+t^{Q}_{\mathrm{H}\to\mathrm{C}}+t^{\mathcal{T}}_{\mathrm{C}}}{\text{d}}t\langle\dot{\Sigma}_{t}\rangle+\int_{t^{\mathcal{T}}_{\mathrm{H}}+t^{Q}_{\mathrm{H}\to\mathrm{C}}+t^{\mathcal{T}}_{\mathrm{C}}}^{t_{\mathrm{f}}}{\text{d}}t\langle\dot{\Sigma}_{t}\rangle\\ =\int_{0}^{t_{\mathrm{f}}}{\text{d}}t\langle\dot{\Sigma}_{t}\rangle\geq 0.$ (S34)
+We proceed to compute the engine’s efficiency, $\langle-W_{\mathrm{cycle}}\rangle/\langle Q_{\mathrm{in}}\rangle$, where $\langle-W_{\mathrm{cycle}}\rangle\equiv-\int_{0}^{t_{\mathrm{f}}}{\text{d}}t\langle\dot{W}_{t}\rangle$ is the total work extracted from the system during one cycle.
+Since we returned to $p(\mathbf{x},t_{\mathrm{f}})=p(\mathbf{x},0)$ and $\lambda^{Q}_{\mathrm{C}\to\mathrm{H}}(t_{\mathrm{f}})=\lambda^{\mathcal{T}}_{\mathrm{H}}(0)$, the mean energy has not changed during one cycle, $\int_{0}^{t_{\mathrm{f}}}{\text{d}}t\langle\dot{H}_{t}\rangle=0$.
+Therefore, according to the first law of thermodynamics [33, 34] [Eq.
+(7) of the main text], the extracted work is $\langle-W_{\mathrm{cycle}}\rangle=\langle Q_{\mathrm{in}}\rangle+\langle Q_{\mathrm{out}}\rangle$.
+Now, since $\mathcal{T}$ is constant for the first and third steps and no heat is involved in the second and fourth steps, and according to Eq.
+(S31) for one cycle we have
+$\frac{\langle Q_{\mathrm{in}}\rangle}{k_{\mathrm{B}}\mathcal{T}_{\mathrm{H}}}+\frac{\langle Q_{\mathrm{out}}\rangle}{k_{\mathrm{B}}\mathcal{T}_{\mathrm{C}}}=-\int_{0}^{t_{\mathrm{f}}}{\text{d}}t\frac{\langle\dot{R}_{t}\rangle}{k_{\mathrm{B}}\mathcal{T}(t)}=-\langle\Sigma_{\mathrm{cycle}}\rangle.$ (S35)
+Thus, we arrive at
+$\frac{\langle-W_{\mathrm{cycle}}\rangle}{\langle Q_{\mathrm{in}}\rangle}=1+\frac{\langle Q_{\mathrm{out}}\rangle}{\langle Q_{\mathrm{in}}\rangle}=1-\frac{{\mathcal{T}}_{\mathrm{C}}}{{\mathcal{T}}_{\mathrm{H}}}-\frac{k_{\mathrm{B}}\mathcal{T}_{\mathrm{C}}\langle\Sigma_{\mathrm{cycle}}\rangle}{\langle Q_{\mathrm{in}}\rangle}.$ (S36)
+Unless one has an estimate of the heat dissipation (informatic entropy production) along a cycle, inserting the bound $\langle\Sigma_{\mathrm{cycle}}\rangle\geq 0$ makes this expression coincide with the Carnot bound for the efficiency of a heat engine, $\langle-W_{\mathrm{cycle}}\rangle/\langle Q_{\mathrm{in}}\rangle\leq 1-{\mathcal{T}}_{\mathrm{C}}/{\mathcal{T}}_{\mathrm{H}}$.
+If the processes were near equilibrium throughout, then $\langle\Sigma_{\mathrm{cycle}}\rangle=0$, and we exactly reach Carnot efficiency.
+Similar to the discussion following our version of the Clausius inequality, it remains to be explored if the temperatures in Eq.
+(S36) are more than just obscure integrated properties of the system which, through a meticulous choice of protocols, were kept constant.
+After all, the temperatures that appear in the classical Carnot bound are those of the two baths, such that the temperature of the system must not coincide with the ones of the baths during irreversible processes.
+SIII.3 Changes in free energy
+It is well known that nonequilibrium steady-states do not necessarily obey a minimum free energy principle [9, 41].
+However, for isothermal transformations we are able to define a free energy that leads to two more relations.
+We begin, once again, from some $p(\mathbf{x},0)$ and perform a protocol $\Lambda(t)=\lambda^{\mathcal{T}}(t)$ which maintains a constant temperature $\mathcal{T}$.
+Define the nonequilibrium Helmholtz free energy $A_{t}=H_{t}-k_{\mathrm{B}}\mathcal{T}S_{t}$.
+Assuming the temperature is indeed constant, $\dot{A}_{t}=\dot{H}_{t}-k_{\mathrm{B}}\mathcal{T}\dot{S}_{t}$.
+Upon inserting Eq.
+(S25) and identifying the work $\dot{W}_{t}=\dot{H}_{t}-\dot{Q}_{t}$, we find
+$\dot{W}_{t}=\dot{A}_{t}+\dot{R}_{t}.$ (S37)
+Taking the ensemble average, we find a bound on the extractable work in noncyclic isothermal processes,
+$\langle-\dot{W}_{t}\rangle=-(\langle\dot{A}_{t}\rangle+\langle\dot{R}_{t}\rangle)\leq-\langle\dot{A}_{t}\rangle,$ (S38)
+which is analogous to the classical relation between extractable work and the change of free energy.
+As mentioned in the main text, since the informatic heat and calorimetric heat are only equal on average and differ otherwise, the Jarzynski equality [37, 38] (that could have reshaped Eq.
+(S38) into an equality) is invalid.
+Lastly, we investigate isothermal processes which occur without external interventions, that is, $\dot{W}_{t}=0$. (This means that $\mathbf{F}=-{\boldsymbol{\nabla}}H$.) We arrive at
+$\langle\dot{A}_{t}\rangle=-\langle\dot{R}_{t}\rangle\leq 0.$ (S39)
+This means that the generalized free energy, like its classical counterpart, keeps decreasing until it reaches steady state in spontaneous isothermal processes.
+This proposes a variational principle for steady-states that result from spontaneous isothermal evolution.
+SIV Example: Steady states with ambiguous generalized temperature
+In this section we discuss a peculiarity related to the nonequilibrium temperature introduced in the main text.
+With a valid Einstein relation $\boldsymbol{\mu}=\mathbf{D}/(k_{\mathrm{B}}T)$ and time-independent conservative forces $\mathbf{F}(\mathbf{x})=-{\boldsymbol{\nabla}}H(\mathbf{x})$, the steady-state distribution of the Fokker-Planck equation
+$\frac{\partial p(\mathbf{x},t)}{\partial t}=-{\boldsymbol{\nabla}}\cdot\mathbf{J}(\mathbf{x},t),\qquad\mathbf{J}(\mathbf{x},t)=-\mathbf{D}(\mathbf{x},t)\cdot\left[\frac{{\boldsymbol{\nabla}}H(\mathbf{x})}{k_{\mathrm{B}}T}p(\mathbf{x},t)+{\boldsymbol{\nabla}}p(\mathbf{x},t)\right]$ (S40)
+is clearly the Boltzmann distribution $p(\mathbf{x},t\to\infty)\sim e^{-H(\mathbf{x})/(k_{\mathrm{B}}T)}$ with the Einstein $T$.
+Now consider the peculiar forcing considered in Sec.
+SII.1, where $\mathbf{F}(\mathbf{x},t)=-\boldsymbol{\mu}^{-1}(\mathbf{x},t)\cdot\mathbf{D}(\mathbf{x},t)\cdot{\boldsymbol{\nabla}}\Phi(\mathbf{x})$ with a scalar function $\Phi(\mathbf{x})$.
+The resulting Fokker-Planck equation,
+$\frac{\partial p(\mathbf{x},t)}{\partial t}=-{\boldsymbol{\nabla}}\cdot\mathbf{J}(\mathbf{x},t),\qquad\mathbf{J}(\mathbf{x},t)=-\mathbf{D}(\mathbf{x},t)\cdot\left[p(\mathbf{x},t){\boldsymbol{\nabla}}\Phi(\mathbf{x})+{\boldsymbol{\nabla}}p(\mathbf{x},t)\right],$ (S41)
+also has a Boltzmann-like steady-state, $p(\mathbf{x},t\to\infty)\sim e^{-\Phi(\mathbf{x})}$, although no temperature appears in the exponent.
+Moreover, $\mathbf{J}(\mathbf{x},t\to\infty)\to\mathbf{0}$, meaning that the proposed nonequilibrium temperature [Eq.
+(15) of the main text] has a ‘$/0$ 0’ ambiguity.
+In an attempt to mend this peculiarity, we consider a two-dimensional system approaching this flux-free steady state.
+Assume that it has constant and diagonal $\boldsymbol{\mu}=\mathrm{diag}(\mu_{1},\mu_{2})$ and $\mathbf{D}=\mathrm{diag}(D_{1}=k_{\mathrm{B}}T_{1}\mu_{1},D_{2}=k_{\mathrm{B}}T_{2}\mu_{2})$.
+For this simplified model of two “particles”, an effective temperature was identified in Ref.
+[69] for the separation vector among them.
+The nonequilibrium temperature we propose here has the following weighted average,
+[…]
+[…]

### v3

@@
 We relate the breaking of the correspondence between entropy production and heat dissipation to departure from the fluctuation-dissipation theorem.
-We propose a temperature-like variable which restores this correspondence and gives rise to a generalized second law of thermodynamics, whereby the dissipated heat is necessarily nonnegative and vanishes at equilibrium.
+We propose a temperaturelike variable that restores this correspondence and gives rise to a generalized second law of thermodynamics, whereby the dissipated heat is necessarily non-negative and vanishes at equilibrium.
 The Clausius inequality, Carnot maximum efficiency theorem, and relation between the extractable work and the change of free energy are recovered as well.
@@
 Among the complications associated with nonequilibrium systems is the inadequacy of temperature and other variables as thermodynamic state functions [7, 8, 9, 10, 11, 12].
-One of the consequences of thermodynamic equilibrium, where the system’s configurations are Boltzmann-distributed, is the Einstein relation (ER) [13, 14],
+One of the consequences of thermodynamic equilibrium, where the system’s configurations are Boltzmann distributed, is the Einstein relation (ER) [13, 14],
 $\mathbf{D}=k_{\mathrm{B}}T\boldsymbol{\mu}\,.$ (1)
@@
 In general they may be space-dependent matrices, reflecting translational and rotational symmetry breaking.
-At equilibrium the effects of fluctuations and dissipation exactly balance [14], leading to Eq.
+At equilibrium with a bath of temperature $T$, the effects of fluctuations and dissipation exactly balance [14], leading to Eq.
 (1).
-However, in cases where forces are not conservative [8, 15], or if fluctuations arise from coarse-graining of fast athermal or active forces [16, 17, 18, 19, 20, 10, 21, 22, 23, 24], there should not be a generic relation between fluctuations (diffusivity) and dissipation (mobility).
-Thus, more often than not, in far-from-equilibrium cases such as active biopolymer networks [25, 26, 27], bacterial swarms [28, 29], and driven chemical reactions [30, 5], the ER does not hold.
-In cases where $\mathbf{D}$ is proportional to $\boldsymbol{\mu}$, one can use Eq.
-(1) to define an effective temperature [7, 10, 31, 19, 18].
-However, in the general case, these two position-dependent tensors are not related by a constant uniform scalar, and Eq.
-(1) is violated.
+Systems away from equilibrium, such as supercooled liquids [15] and systems under external forces [8], typically do not satisfy Eq.
+(1) with the ambient temperature $T$.
+In such cases Eq.
+(1) has been used to define an effective temperature [7, 10, 16, 17, 18].
+However, if the forces are not conservative [8, 19], or when fluctuations arise from coarse graining of fast athermal or active forces [20, 21, 18, 17, 22, 10, 23, 24, 25, 26], there is generally no relation between the two position-dependent tensors $\mathbf{D}$ and $\boldsymbol{\mu}$.
+Thus, in far-from-equilibrium cases such as active biopolymer networks [27, 28, 29], bacterial swarms [30, 31], and driven chemical reactions [32, 5], the ER is invalid, as no scalar $T$ exists to satisfy Eq.
+(1).
+In this Letter, we propose a temperaturelike variable for systems exhibiting this severe breakdown of the Einstein relation.
 One of the theories addressing nonequilibrium systems is stochastic thermodynamics [33].
-By considering Langevin equations or discrete Markov chains it establishes trajectory-wise energy conservation (leading to the first law of thermodynamics) [33, 34, 35] and gives a wealth of fluctuation relations (FRs) [36], such as the Jarzynski and Crooks relations [37, 38, 39].
+By considering Langevin equations or discrete Markov chains, it establishes trajectorywise energy conservation (leading to the first law of thermodynamics) [34, 35, 36] and gives a wealth of fluctuation relations (FRs) [37], such as the Jarzynski and Crooks relations [38, 39, 40].
 The latter also yields the second law of thermodynamics for the average entropy changes and heat increments.
-Yet, implicitly, all the aforementioned FRs and extended laws of thermodynamics have relied on the existence of a temperature, defined through the ER, to connect microscopic statistics to thermodynamic energy, mechanical and chemical work, and calorimetric heat.
+Yet, implicitly, all the aforementioned FRs and extended second law of thermodynamics have relied on the existence of a temperature, defined through the ER, to connect microscopic statistics to thermodynamic energy, mechanical and chemical work, and calorimetric heat.
 Several definitions of a nonequilibrium temperature were proposed based on generalizations of the fluctuation-dissipation theorem (FDT) and FRs in active matter [10, 22, 41, 42], systems with memory [20, 21], and anomalous diffusion [43, 44].
 Most of these require a thermal bath that obeys the ER.
-In this Letter, we adopt an alternative approach, which is not derived from extentions of FRs or FDT, to identify a temperature-like variable directly, such that the classical-thermodynamic results are recovered.
+In this Letter, we adopt an alternative approach, which is not derived from extensions of FRs or FDT, to identify a temperaturelike variable directly, such that the classical-thermodynamic results are recovered.
 In some cases, the ER is required to keep the extended FRs connected to physical measurables.
@@
 Otherwise, they demand knowledge of the complete microstates’ steady-state distribution, which is usually inaccessible or requires simplistic models [46].
-The temperature-like variable proposed here, arising directly from thermodynamics, does not rely on those assumptions.
+The temperaturelike variable proposed here, arising directly from thermodynamics, does not rely on those assumptions.
 We have two aims: (i) to investigate the fate of existing nonequilibrium FRs and extensions to the second law of thermodynamics once ER is violated; and (ii) to formulate a generalized, ER-independent temperature and derive the resulting alternative thermodynamic identities.
@@
 We denote the total ($N$-component) force acting on the particles by $\mathbf{F}=\mathbf{F}_{\mathrm{ext}}-\boldsymbol{\nabla}H$, where $H(\mathbf{x},\lambda)$ is the system’s Hamiltonian.
-The system is driven out of equilibrium due to the nonconservative force $\mathbf{F}_{\mathrm{ext}}$ and/or violation of ER [Eq.
+The system is driven out of equilibrium due to the nonconservative force $\mathbf{F}_{\mathrm{ext}}$ [47] and/or violation of ER [Eq.
 (1)].
@@
 $+\mathbf{b}(\mathbf{X}_{t},\Lambda_{t})\cdot\boldsymbol{\eta}_{t}\,,$
-where $\mathbf{b}$ is the noise magnitute satisfying $\mathbf{D}=\mathbf{b}\cdot\mathbf{b}^{\mathrm{T}}/2$, $\boldsymbol{\eta}_{t}$ is the normalized thermal noise [$\langle\boldsymbol{\eta}_{t}\rangle=\mathbf{0}$, $\langle\boldsymbol{\eta}_{t}\boldsymbol{\eta}_{t^{\prime}}\rangle=\mathbf{I}\delta(t-t^{\prime})$ where $\mathbf{I}$ is the identity matrix], and the equation is written in the Itô convention [46].
+where $\mathbf{b}$ is the noise magnitude satisfying $\mathbf{D}=\mathbf{b}\cdot\mathbf{b}^{\mathrm{T}}/2$, $\boldsymbol{\eta}_{t}$ is the normalized thermal noise [$\langle\boldsymbol{\eta}_{t}\rangle=\mathbf{0}$, $\langle\boldsymbol{\eta}_{t}\boldsymbol{\eta}_{t^{\prime}}\rangle=\mathbf{I}\delta(t-t^{\prime})$, where $\mathbf{I}$ is the identity matrix], and the equation is written in the Itô convention [48].
 The “spurious drift” term ${\boldsymbol{\nabla}}\cdot\mathbf{D}$ ensures that Boltzmann distribution is obtained at equilibrium, i.e., in systems with a well-defined temperature, Eq.
@@
 where $\mathbf{J}$ is the probability flux.
-We follow Sekimoto’s approach [33, 34] to identify the internal energy, work, and heat for trajectories in generic Langevin systems. The time-derivative of the Hamiltonian is [46]
+We follow Sekimoto’s approach [34, 35] to identify the internal energy, work, and heat for trajectories in generic Langevin systems. The time derivative of the Hamiltonian is [48]
 $\dot{H}_{t}={\boldsymbol{\nabla}}H\circ\dot{\mathbf{X}}_{t}+\frac{\partial H}{\partial\lambda}\dot{\Lambda}_{t}\,,$ (5)
@@
 Combining the stochastic chain rule [48] and Eqs.
-(2)–(4) yields the time-derivative of $S_{t}=S(\mathbf{X}_{t},t)$,
+(2)–(4) yields the time derivative of $S_{t}=S(\mathbf{X}_{t},t)$,
 $\dot{S}_{t}={\boldsymbol{\nabla}}\cdot\left(\frac{\mathbf{J}}{p}\right)-\frac{1}{p}{\boldsymbol{\nabla}}\cdot(\mathbf{D}\cdot{\boldsymbol{\nabla}}p)-[({\boldsymbol{\nabla}}\ln p)\cdot\mathbf{b}]\cdot\boldsymbol{\eta}_{t}\,,$ (9)
@@
 It quantifies how much extra calorimetric heat $\dot{Q}_{t}$ is wasted in an irreversible process.
-According to the second law of thermodynamics, the average thermodynamic dissipation $\langle\dot{R}_{t}\rangle$ is nonnegative [56],
+According to the second law of thermodynamics, the average thermodynamic dissipation $\langle\dot{R}_{t}\rangle$ is non-negative [58],
 $\dot{R}_{t}\equiv\dot{S}_{t}-\frac{\dot{Q}_{t}}{k_{\mathrm{B}}T}\,,\qquad\langle\dot{R}_{t}\rangle\geq 0\,.$ (10)
@@
 Consider a discretized trajectory drawn from Eq.
-(2), $(\mathbf{X}_{0},\dots,\mathbf{X}_{K\Delta t})=(\mathbf{x}_{0},\dots,\mathbf{x}_{K})$, with arbitrarily-small timestep $\Delta t$.
+(2), $(\mathbf{X}_{0},\dots,\mathbf{X}_{K\Delta t})=(\mathbf{x}_{0},\dots,\mathbf{x}_{K})$, with arbitrarily small time step $\Delta t$.
 The probability distribution of obtaining this trajectory under the protocol sequence $(\lambda_{0},\lambda_{1},\ldots,\lambda_{K})$, $\overrightarrow{\mathcal{P}}\equiv\Pr(\mathbf{x}_{0},\mathbf{x}_{1},\ldots,\mathbf{x}_{K})$, is compared with the probability distribution of obtaining the reverted trajectory, $\overleftarrow{\mathcal{P}}\equiv\Pr(\mathbf{x}_{K},\mathbf{x}_{K-1},\ldots,\mathbf{x}_{0})$, under the reverted protocol $(\lambda_{K},\lambda_{K-1},\ldots,\lambda_{0})$.
@@
 In Eq.
-(12) we identify the time-derivative of the entropy, $\dot{S}_{t}=\ln[p(\mathbf{x}_{k+1},(k+1)\Delta t)/p(\mathbf{x}_{k},k\Delta t)]/\Delta t$, given by Eq.
+(12) we identify the time derivative of the entropy, $\dot{S}_{t}=\ln[p(\mathbf{x}_{k+1},(k+1)\Delta t)/p(\mathbf{x}_{k},k\Delta t)]/\Delta t$, given by Eq.
 (9).
@@
 Therefore, if $\mathbf{D}$ and $\boldsymbol{\mu}$ are not related by a scalar, the informatic $\dot{\Sigma}_{t}$ is no longer related to the measurable $\dot{R}_{t}$, even on average.
-Consequently, the nonnegativity of $\langle\dot{R}_{t}\rangle$ (i.e., the second law) can no longer be established based on trajectory statistics.
+Consequently, the non-negativity of $\langle\dot{R}_{t}\rangle$ (i.e., the second law) can no longer be established based on trajectory statistics.
 Below we propose a definition of a nonequilibrium temperature that mends this discrepancy.
@@
 We now impose the following physical conditions on Eq.
-(14): (i) To satisfy the second law, $\langle\dot{R}_{t}\rangle$ should be a nonnegative functional of $\mathbf{J}$.
+(14): (i) To satisfy the second law, $\langle\dot{R}_{t}\rangle$ should be a non-negative functional of $\mathbf{J}$.
 (ii) To agree with Onsager’s theory [63, 64, 65], the right-hand side should reduce to a quadratic form of $\mathbf{J}$ near equilibrium and turn the second-law inequality into an equality for reversible processes ($\mathbf{J}\to\mathbf{0}$).
@@
 $k_{\mathrm{B}}\mathcal{T}=\frac{\int{\text{d}}\mathbf{x}\,\mathbf{J}\cdot\mathbf{F}}{\int{\text{d}}\mathbf{x}\,\mathbf{J}\cdot\mathbf{D}^{-1}\cdot(\boldsymbol{\mu}\cdot\mathbf{F})}\,.$ (15)
-We have thus recovered the second law of thermodynamics.
+We have thus recovered the second law of thermodynamics for general Markovian overdamped Langevin dynamics (Eq. (2)).
 Equation (15) is our central result.
@@
 Otherwise, the generalized temperature is an intricate quantity.
+Computing it directly from Eq.
+(15) is a challenge.
+Below we propose how to estimate it from experiments (see, e.g., Eq.
+(18)).
 This is expected for nonequilibrium systems with arbitrarily complex $\mathbf{J}$, $\mathbf{D}$, $\boldsymbol{\mu}$, and $\mathbf{F}$.
@@
 $+[(\boldsymbol{\mu}\cdot\mathbf{F})\cdot\mathbf{D}^{-1}\cdot\mathbf{b}]\cdot\boldsymbol{\eta}_{t}\,,$
-which, combined with Eq. (9) and the relation $\dot{\Sigma}_{t}=\dot{S}_{t}+\dot{\Omega}_{t}$, gives the IEP rate as
+which, combined with Eq. (9) and the relation $\dot{\Sigma}_{t}=\dot{S}_{t}+\dot{\Omega}_{t}$, gives the IEP rate as (cf. Ref. [67])
 $\dot{\Sigma}_{t}=\frac{\mathbf{J}}{p}\cdot\mathbf{D}^{-1}\cdot\frac{\mathbf{J}}{p}+\frac{2}{p}{\boldsymbol{\nabla}}\cdot\mathbf{J}+\left(\frac{\mathbf{J}}{p}\cdot\mathbf{D}^{-1}\cdot\mathbf{b}\right)\cdot\boldsymbol{\eta}_{t}\,.$ (17)
@@
 For example, the variances of the former are different, $\langle({\text{d}}\Omega_{t})^{2}\rangle_{\mathrm{c}}=2{\text{d}}t\int(\boldsymbol{\mu}\cdot\mathbf{F})\cdot\mathbf{D}\cdot(\boldsymbol{\mu}\cdot\mathbf{F})p{\text{d}}\mathbf{x}$, whereas $\langle({\text{d}}Q_{t})^{2}\rangle_{\mathrm{c}}=2{\text{d}}t\int\mathbf{F}\cdot\mathbf{D}\cdot\mathbf{F}p{\text{d}}\mathbf{x}\neq(k_{\mathrm{B}}\mathcal{T})^{2}\langle({\text{d}}\Omega_{t})^{2}\rangle_{\mathrm{c}}$.
-(The subscript ${\mathrm{c}}$ denotes a cumulant.) Since the Jarzynski equality [37, 38] and the Crooks fluctuation theorem [39] require statistics beyond the mean, they are not restored by the generalized $\mathcal{T}$.
+(The subscript ${\mathrm{c}}$ denotes a cumulant.) Since the Jarzynski equality [38, 39] and Crooks fluctuation theorem [40] require statistics beyond the mean, they are not restored by the generalized $\mathcal{T}$.
 Thus, without the ER, the various FRs [37] remain related only to informatic quantities (the IEP and IH) [45] and not to thermodynamics (dissipation and heat).
@@
 Nevertheless, generalized versions of the FDT and the Harada-Sasa relation can be derived in the absence of the ER.
-Here we outline these results; see the Supplementary Material (SM) for details [67].
+Here we outline these results; see the Supplemental Material for details [70].
 Instead of considering the response to an impulsive linear potential gradient as was done in Ref.
@@
 If, instead of $\mathbf{F}$, we consider the observable $\mathbf{D}^{-1}\cdot\boldsymbol{\mu}\cdot\mathbf{F}$, we obtain a similar relation for the IH,
-$\langle\dot{\Omega}_{s}\rangle=\lim_{t\to s^{+}}\left(\frac{{\text{d}}}{{\text{d}}s}\langle[(\boldsymbol{\mu}\cdot\mathbf{F})\cdot\mathbf{D}](\mathbf{X}_{t},t)\cdot\mathbf{X}_{s}\rangle\right.\\ \left.+\frac{\delta}{\delta\mathbf{h}(s)}\cdot\langle[(\boldsymbol{\mu}\cdot\mathbf{F})\cdot\mathbf{D}](\mathbf{X}_{t},t)\rangle\right)\,.$ (21)
+$\langle\dot{\Omega}_{s}\rangle=\lim_{t\to s^{+}}\left(\frac{{\text{d}}}{{\text{d}}s}\langle[(\boldsymbol{\mu}\cdot\mathbf{F})\cdot\mathbf{D}^{-1}](\mathbf{X}_{t},t)\cdot\mathbf{X}_{s}\rangle\right.\\ \left.+\frac{\delta}{\delta\mathbf{h}(s)}\cdot\langle[(\boldsymbol{\mu}\cdot\mathbf{F})\cdot\mathbf{D}^{-1}](\mathbf{X}_{t},t)\rangle\right)\,.$ (21)
 Thus the IH and thermodynamic heat are related to violations of a generalized FDT for different observables.
@@
 We now revisit additional classical-thermodynamic results demonstrating the relevance of the generalized temperature $\mathcal{T}$.
-More details regarding these consequences are found in the SM [67].
+More details regarding these consequences are found in the Supplemental Material [70].
 (i) Consider a process that starts and ends at the same protocol value, $\Lambda_{t_{\mathrm{f}}}=\Lambda_{0}$, and then allowed to relax to steady state.
@@
 [71] in the context of nonequilibrium engines.) (iii) Define the nonequilibrium free energy, $A_{t}=H_{t}-k_{\mathrm{B}}\mathcal{T}S_{t}$.
-It is easy to show that it bounds the extractable work in “isothermal” (constant $\mathcal{T}$) transformations, $\langle-\dot{W}_{t}\rangle=-(\langle\dot{A}_{t}\rangle+k_{\mathrm{B}}\mathcal{T}\langle\dot{\Sigma}_{t}\rangle)\leq-\langle\dot{A}_{t}\rangle$ [67].
+It is easy to show that it bounds the extractable work in isothermal (constant $\mathcal{T}>0$) transformations, $\langle-\dot{W}_{t}\rangle=-(\langle\dot{A}_{t}\rangle+k_{\mathrm{B}}\mathcal{T}\langle\dot{\Sigma}_{t}\rangle)\leq-\langle\dot{A}_{t}\rangle$ [70].
 An interesting consequence of Eq.
@@
 These models may be relevant, for instance, for mixtures of passive and active particles [74, 75, 76, 77, 78].
-They yield a temperature-like scalar which is a weighted average of the thermostats’ temperatures.
+They yield a temperaturelike scalar that is a weighted average of the thermostats’ temperatures.
 Considering a diagonal but nonscalar $\mathbf{D}^{-1}\cdot\boldsymbol{\mu}$, with different inverse temperatures as its eigenvalues, we may view Eq.
 (15) as a generalization of such a weighted average.
-These systems may attain $\mathbf{J}=\mathbf{0}$, which results in a ‘$/0$ 0’ ambiguity for $\mathcal{T}$.
+These systems may attain $\mathbf{J}=\mathbf{0}$, which results in a “$/0$ 0” ambiguity for $\mathcal{T}$.
 However, perturbing different global parameters amounts to taking a perturbation $\mathbf{J}$ along different axes, thus picking up different eigenvalues (inverse temperatures) of $\mathbf{D}^{-1}\cdot\boldsymbol{\mu}$.
-We show a simplistic example in the SM [67].
+We show a simplistic example in the Supplemental Material [70].
 Conclusion.—We have analyzed a generic nonequilibrium Langevin system [Eq.
-(2)] which violates the ER [Eq.
+(2)] that violates the ER [Eq.
 (1)].
@@
 To restore the information-thermodynamics correspondence we have imposed the second law of thermodynamics directly and identified a generalized temperature $\mathcal{T}$ [Eq.
-(15)], which gives all the thermodynamically required properties: nonnegative dissipation, zero dissipation at equilibrium, linear thermodynamic stability, and the temperature being a macroscopic (integrated) property.
-The generalized temperature is model-independent — with a calorimetric measurement of the heat and path-sampling for the IH, one may find the system’s $\mathcal{T}$ with no required knowledge of the underlying dynamics.
+(15)], which gives all the thermodynamically required properties: non-negative dissipation, zero dissipation at equilibrium, linear thermodynamic stability, and the temperature being a macroscopic (integrated) property.
+The generalized temperature is model-independent—with a calorimetric measurement of the heat and path sampling for the IH, one may find the system’s $\mathcal{T}$ with no required knowledge of the underlying dynamics.
 In particular, the variety of tools available for the estimation of IEP from trajectories (e.g., sampling marginal distributions [79, 80, 60] or the thermodynamic uncertainty relations [81, 82]) will be useful, since in our construction the average IEP remains connected to the average thermodynamic dissipation.
 The generalized temperature does not restore the various FRs [38, 40, 37, 68], which in the absence of the ER no longer dictate the statistics of thermodynamic properties [45].
-The three conditions which have been used to find $\mathcal{T}$, Eq.
+The three conditions that have been used to find $\mathcal{T}$, Eq.
 (15), do not uniquely determine it.
@@
 Acknowledgements.
-We greatly benefited from discussions with David Dean and Yael Roichman. T.M. acknowledges funding from the Israel Science Foundation (Grant No. 1356/22).
+We greatly benefited from discussions with David Dean and Yael Roichman.
+T.M.
+acknowledges funding from the Israel Science Foundation (Grant No.
+1356/22).
+G.A.
+and H.D.
+acknowledge funding from the Israel Science Foundation (Grant No.
+1611/24).
 Supplementary Material: Second Law of Thermodynamics without Einstein Relation
@@
   • For reversible (near-equilibrium) processes, the Clausius inequality is an equality.
-Here, even if the Einstein relation does not apply, we find that at flux-free steady state, $\mathbf{J}=\mathbf{0}$, the generalized Clausius inequality becomes an equality.
+Here, even if the Einstein relation does not apply, we find that at flux-free steady state, $\mathbf{J}=\mathbf{0}$, the generalized Clausius ineqsuality becomes an equality.
 Moreover, it indicates that the flux-free state is stable (that is, for a small departure from equilibrium $\delta\mathbf{J}$, the dissipation is quadratic in it and hence much smaller).